### v1

# Some remarks on the exponential separation and dimension preserving approximation for sets and measures

Abstract.
In the dimension theory of sets and measures, a recent breakthrough happened due to Hochman, who introduced the exponential separation condition and proved the Hausdorff dimension result for invariant sets and measures generated by similarities on the real line. Following this groundbreaking work, we make a modest contribution by weakening the condition. We also define some sets in the class of all nonempty compact sets using the Assouad and Hausdorff dimensions and subsets of measures in the Borel probability measures using the $L^{q}$ dimension and prove their density in the respective spaces.
Key words and phrases:
Separation condition, self-similar sets, invariant measures, Assouad dimension, $L^{q}$ dimension, Hausdorff dimension
2020 Mathematics Subject Classification:
Primary 28A80; Secondary 28A78
1. Introduction
The dimension estimation of a fractal set and associated measure constructed using an iterated function system (IFS) (see [20]) has attracted a lot of researchers over the last century. Depending on the properties of the contractive transformations, such as similarity (also called similitude), affine, or conformal in $\mathbb{R}^{m},$ the fractal set and the measure are characterized as self-similar, self-affine, and self-conformal, respectively. Several dimensions are introduced in the literature, as an example Hausdorff dimension $(\dim_{\mathcal{H}})$ [11, 22], Assouad dimensions $(\dim_{\mathcal{A}})$ [14], $L^{q}$ dimension $(\dim_{L^{q}})$ [24, 30] for the invariant set and measure. Initially, we mention recent progress being made in dimension estimation. Let $A$ and $\mu$ be the fractal set and measure associated to some IFS on $\mathbb{R}^{m}$. The trivial bound is obtained as
$\dim_{\mathcal{H}}A\leq\min\{m,\dim_{S}A\},$
$\dim_{\mathcal{H}}\mu\leq\min\{m,\dim_{S}\mu\},$
[…]
  • Under OSC, $\dim_{\mathcal{H}}A=\min\{m,\dim_{S}A\},$ see [20, 23].
  • IFS with bi-Lipschitz mappings $(s_{i}\|x-y\|_{2}\leq\|f_{i}(x)-f_{i}(y)\|_{2}\leq t_{i}\|x-y\|_{2})$ on $\mathbb{R}^{m}$ and SSC, we have $s\leq\dim_{\mathcal{H}}A\leq t,$ where $s$ and $t$ satisfies $\sum_{i=1}^{N}s_{i}^{s}=1$ and $\sum_{i=1}^{N}t_{i}^{t}=1,$ respectively, see [11, Proposition 9.6-9.7]. It is noted that the upper bound holds without any separation condition, but the lower bound does not hold even if we assume OSC. Additionally, it does not serve as a bound for the Assouad dimension. For example, visit [14, Section 8.5].
  • Total overlap implies $\dim_{\mathcal{H}}A<\min\{m,\dim_{S}A\}.$
  • Generally, $\dim_{\mathcal{H}}A\leq\dim_{\mathcal{A}}A,$ but if $A$ is Ahlfors regular, then $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A,$ see [14, Theorem 6.4.1]. IFS with similitudes on $\mathbb{R},$ either $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A$ or $\dim_{\mathcal{A}}A=1.$
[…]
  • In $\mathbb{R}$ with similarities and ESC, Shmerkin computed $\dim_{L^{q}}\mu=\min\{1,\tilde{\tau}(q)/(q-1)\},$ see [30, Theorem 6.6].
  • IFS with similarity and OSC implies $\dim_{\mathcal{H}}\mu=\min\{m,\dim_{S}\mu\}$, see [10]. In this case, under SSC, Fraser and Howroyd computed $\dim_{\mathcal{A}}\mu,$ see [14, Theorem 7.4.1].
It is well known that the Hausdorff dimension decreases under Lipschitz mapping ([11, Corollary 2.4]), but this is not the case with the Assouad dimension ([14, Corollary 7.2.2]). Bedford [6] in his PhD thesis, studied the box dimension of self-similar sets. Recent work in the dimension estimation of self-similar sets can be explored in [35]. Initially, a special class of self-affine sets known as self-affine carpets, such as Bedford-McMullen carpets, is studied; for details, see [15]. Several recent works in the direction of estimation of the Hausdorff dimension of self-affine sets and measures can be found in [4, 12, 19, 21, 31].
The primary theorems in this article include the following.
  • The weaker version of the Hochman result [18, Theorem 1.4] is shown in Theorem 3.15.
  • The modified ESC is being defined using the Hausdorff metric and compared with the existing ESC, along with some interesting remarks.
  • Several dense subsets are explored that preserve the Assouad dimension and $L^{q}$ dimension.
Finally, the identified research gap and potential future comments are outlined in the last Section 4.
2. Preliminaries
Let $\mathcal{K}(\mathbb{R}^{m})$ denote the collection of all nonempty compact subsets of $\mathbb{R}^{m}$ in the Euclidean norm $\|.\|_{2}.$ The Hausdorff distance $\mathfrak{H}$ between any $A,B\in\mathcal{K}(\mathbb{R}^{m})$ is elaborated as
$\mathfrak{H}(A,B):=\inf\{\epsilon>0:A\subset B_{\epsilon}\text{ and }B\subset A_{\epsilon}\},$
where $A_{\epsilon}$ denotes the $\epsilon$-neighbourhood of $A.$ It is also redefined as
$\mathfrak{H}(A,B):=\max\Big{\{}\sup_{a\in A}\inf_{b\in B}\|a-b\|_{2},\sup_{b\in B}\inf_{a\in A}\|a-b\|_{2}\Big{\}}.$
Definition 2.1.
[11] Let $E$ be a subset of $\mathbb{R}^{m}$ and $r$ be any non-negative number. Then, the $r$-dimensional Hausdorff measure of $E$ is defined as
$\mathcal{H}^{r}(E)=\lim_{\delta\to 0^{+}}\mathcal{H}_{\delta}^{r}(E),$
where $\mathcal{H}_{\delta}^{r}(E)=\inf\{\sum_{i=1}^{\infty}(diam(U_{i}))^{r}:~\{U_{i}\}\text{ is a }\delta\text{-cover of }E\}$ and $diam(U_{i})=\sup_{x,y\in U_{i}}\|x-y\|_{2}.$ The Hausdorff dimension of $E$ is defined as
$\dim_{\mathcal{H}}E=\inf\{r\geq 0:~\mathcal{H}^{r}(E)=0\}.$
The approximation of a compact set that preserves dimension is explored in [36].
Definition 2.2.
[14] Let $E$ be a non-empty subset of $\mathbb{R}^{m}.$ The Assouad dimension of $E$ is defined as
$\dim_{\mathcal{A}}E=\inf\Bigg{\{}\beta\geq 0:\exists\text{ a constant }$ $C>0\text{ such that }\forall~0<r<R\text{ and }x\in E,$
$N_{r}(B(x,R)\cap E)\leq C\Big{(}\frac{R}{r}\Big{)}^{\beta}\Bigg{\}},$
where $B(x,R)=\{y\in\mathbb{R}^{m}:\|y-x\|_{2}\leq R\}$ and $N_{r}(B(x,R)\cap E)$ is the smallest number of open set required for a $r$-cover of $B(x,R)\cap E.$
Example 2.3.
The Assouad dimension of any finite subset of $\mathbb{R}^{m}$ is zero.
Chen [8] et al. explored the accessible values for the Assouad dimension of subsets. We now mention properties of the Assouad dimension in the following theorem, used in the article.
Theorem 2.4.
[14, Lemma 2.4.1, Lemma 2.4.2]
  1. (1)
    The Assouad dimension is monotone, finitely stable, stable under closure, and satisfies the open set property.
  2. (2)
    Let $E$ be a nonempty subset of $\mathbb{R}^{m},$ and $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ be a bi-Lipschitz map. Then, $\dim_{\mathcal{A}}T(E)=\dim_{\mathcal{A}}E.$
Let $\mathcal{P}(\mathbb{R}^{m})$ be the set of probability measures on $\mathbb{R}^{m}$ with the metric defined as
$d_{L}(\mu,\nu)=\sup_{Lip(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\mu-\int_{\mathbb{R}^{m}}fd\nu\Big{|}\Big{\}},$
where $f$ is a Lipschitz function with the Lipschitz constant $Lip(f)$ equal to $1$, see [7, 25]. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, the convolution $\mu*\nu$ is given as
$\mu*\nu(A)=\int_{\mathbb{R}^{m}}\mu(A-x)d\nu(x),\hskip 11.38092pt\mbox{for all Borel set }A\in\mathbb{R}^{m}.$
Equivalently, $\mu*\nu$ is the unique Borel probability measure satisfying
$\int_{\mathbb{R}^{m}}f(x)d(\mu*\nu)dx=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\mu(x)d\nu(y),~\forall~f\in C_{b}(\mathbb{R}^{m}),$
where $C_{b}(\mathbb{R}^{m})$ denotes the set of all continuous and bounded functions on $\mathbb{R}^{m}$. We know that (i) $\mu*\nu(A)=\nu*\mu(A)$ and (ii) for any $x_{0}\in\mathbb{R}^{m},$ we have $\mu*\delta_{x_{0}}(A)=\mu(A-x_{0}).$ Feng et al. [13] studied the convolutions of equicontractive self-similar measures on $\mathbb{R}.$
Definition 2.5.
[30] Let $q\in(1,\infty).$ If $\mu$ is a probability measure on $\mathbb{R}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
$\tau(\mu,q)=\liminf_{k\to\infty}\frac{-\log\sum_{I\in D_{k}}\mu(I)^{q}}{k},$
where $D_{k}=\{[j2^{-k},(j+1)2^{-k}):j\in\mathbb{Z}\}$ is the level-$k$ dyadic partitions of $\mathbb{R}.$ The $L^{q}$ dimension of $\mu$ is given as $\dim_{L^{q}}(\mu)=\frac{\tau(\mu,q)}{q-1}.$
If $\mu$ is a probability measure on $\mathbb{R}^{m}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
$\tau(\mu,q)=\liminf_{k\to\infty}\frac{-\log\sum_{I^{1}\times I^{2}\times\cdots\times I^{m}\in D_{k}^{m}}\mu(I^{1}\times I^{2}\times\cdots\times I^{m})^{q}}{k},$
where $D_{k}^{m}=\{[j2^{-k},(j+1)2^{-k}):j\in\mathbb{Z}\}^{m}$ is the dyadic partitions of $\mathbb{R}^{m}.$ The $L^{q}$ dimension of $\mu$ is given as $\dim_{L^{q}}(\mu)=\frac{\tau(\mu,q)}{q-1}.$
For entropy and $L^{q}$ dimension, researchers consider preferring logarithms on base $2,$ for normalizing the value of dimension. Several recent works on $L^{q}$ dimension can be studied in [2, 24, 30].
Example 2.6.
We now proceed to estimate the $L^{q}$ dimension of some well-known probability measures, to enhance clarity for new readers.
  • Let $x_{0}\in\mathbb{R}$ and $\delta_{x_{0}}$ be the Dirac measure supported on $x_{0}.$ Then,
    $\tau(\delta_{x_{0}},q)=\liminf_{k\rightarrow\infty}\frac{-\log\sum_{I\in D_{k}}(\delta_{x_{0}}(I))^{q}}{k}=\liminf_{k\rightarrow\infty}\frac{-\log(1)^{q}}{k}=0.$
    This gives $\dim_{L^{q}}(\delta_{x_{0}})=0.$
  • Let $c\in\mathbb{R}.$ Then,
    $\tau(c\delta_{x_{0}},q)=\liminf_{k\rightarrow\infty}\frac{-\log(c)^{q}}{k}=0,\text{ and }\dim_{L^{q}}(c\delta_{x_{0}})=0.$
  • Let $x_{1},x_{2},\ldots,x_{n}$ be distinct points in $\mathbb{R}.$ Then,
    $\tau(c_{1}\delta_{x_{1}}+c_{2}\delta_{x_{2}}+\dots+c_{n}\delta_{x_{n}},q)=\liminf_{k\rightarrow\infty}\frac{-\log(c_{1}^{q}+c_{2}^{q}+\ldots+c_{n}^{q})}{k}=0,$
    and $\dim_{L^{q}}(c_{1}\delta_{x_{1}}+c_{2}\delta_{x_{2}}+\dots+c_{n}\delta_{x_{n}})=0.$
  • Let $\mathcal{L}\big{|}_{[0,1]}$ be the normalized Lebesgue measure supported on $[0,1].$ Then,
    $\tau(\mathcal{L}\big{|}_{[0,1]},q)$ $=\liminf_{k\rightarrow\infty}\frac{-\log\sum_{I\in D_{k}}(\mathcal{L}\big{|}_{[0,1]}(I))^{q}}{k}$
    $=\liminf_{k\rightarrow\infty}\frac{-\log(2^{k}-1)2^{-kq}}{k}$
    $=\liminf_{k\rightarrow\infty}\frac{-\log(2^{k}-1)}{k}+\liminf_{k\rightarrow\infty}\frac{-\log 2^{-kq}}{k}$
    $=-\log 2+q\log 2=(q-1)\log 2=q-1.$
    Thus, $\dim_{L^{q}}(\mathcal{L}\big{|}_{[0,1]})=1$.
Let $\Phi$ denote the similarity mapping on $\mathbb{R}^{m},$ then $\Phi(x)=s\mathcal{O}x+d,$ where $s\in(-1,1)$ is the scaling ratio, $\mathcal{O}$ is the $m\times m$ orthogonal matrix, and $d\in\mathbb{R}^{m}$ is the translation vector. Hochman [18] defined the distance between two similarities $\Phi$ and $\Phi_{*}$ in $\mathbb{R}^{d}$ as
$d(\Phi,\Phi_{*}):=|\log s-\log s_{*}|+\|\mathcal{O}-\mathcal{O}_{*}\|+\|d-d_{*}\|_{2},$
where $\|.\|_{2}$ denotes the Euclidean norm in $\mathbb{R}^{m}$ and $\|.\|$ is the operator norm. Consider an IFS that consists of a finite number of similarity mappings on $\mathbb{R}^{n}$, i.e., $\mathcal{I}=\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ and
(2.1) $\Delta_{n}=\inf\{d(h_{i},h_{j}):i,j\in\{1,2,\ldots,N\}^{n},i\neq j\}.$
For $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots N\}^{n},$ we denote $h_{i}=h_{i_{1}}\circ h_{i_{2}}\circ\cdots\circ h_{i_{n}}.$ An IFS $\mathcal{I}$ is said to satisfy the exponential separation condition (ESC) if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ see [19].
We would like to highlight that in some of the articles, the previous definition of ESC is mentioned with all arbitrarily large $n\in\mathbb{N}$ such that $\Delta_{n}>\delta^{n},$ while some others say that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N},$ for example refer to the literature [3, 18, 30]. Observe that if $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ then both are true. In the other case, if $\Delta_{n}>\delta^{n},$ for all $n\geq n_{0},$ then choose $\delta_{*}=\min\{\delta,\Delta_{1},\sqrt{\Delta_{2}}\ldots,\sqrt[n_{0}]{\Delta_{n_{0}}}\}.$ This gives $\Delta_{n}>\delta_{*}^{n},$ for all $n\in\mathbb{N}.$ With reference to Theorem 1.5 of [17], for an IFS on $\mathbb{R}$ consisting of all the similarity mappings with the algebraic parameters, $\Delta_{n}$ is either zero or strictly greater than $\delta^{n},$ for some constant $\delta>0,$ establishing the equivalency among the three. We strongly believe that all three definitions are equivalent in general. Further, note that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N}$ is equivalent to $\Delta_{n}>c\delta^{n}$ for all $n\in\mathbb{N},$ for some constant $c\in\mathbb{R}.$ The direct part is straightforward. For the converse, consider two cases. First $c>1,$ then it is obvious. In the other case, assume $0<c\leq 1.$ Choose $\delta_{*}=c\delta,$ then $\Delta_{n}>c\delta^{n}>(c\delta)^{n}.$ An example was presented by Garsia [16] as $\{\mathbb{R}:~h_{1}(x)=\gamma^{-1}x+1,h_{2}(x)=\gamma^{-1}x-1\},$ where $\gamma$ is the real root of the equation $x^{3}-x^{2}-2=0.$ Hochman [18] shows that this IFS satisfies ESC, but not OSC. Meanwhile, super-exponential condensation means $\lim_{n\rightarrow\infty}\frac{1}{n}\log\Delta_{n}=-\infty.$ Bárány and Käenmäki [3, Theorem 2.1] showed the existence of an uncountable family of homogeneous self-similar sets of $\mathbb{R}$ that do not have total overlaps and satisfy super-exponential condensation. Baker [1] also independently showed the existence of such self-similar sets in $\mathbb{R}$ using the theory of continued fractions. Using the technique of Baker, Chen [9] provided some more examples in $\mathbb{R}.$
3. Main Results
Lemma 3.1.
Let $A$ and $B$ be the matrices of order $k\times k$ and $m\times m,$ respectively. Then,
$\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\Big{\|}\geq\max\{\|A\|,\|B\|\},$
where $\textbf{0}_{1}$ and $\textbf{0}_{2}$ are zero matrices of order $k\times m$ and $m\times k,$ respectively. Further, if $A$ and $B$ are orthogonal, so is $\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}$.
Proof.
Let $x=(x_{1},x_{2},\ldots,x_{k})\in\mathbb{R}^{k}$ and $y=(y_{1},y_{2},\ldots,y_{m})\in\mathbb{R}^{m}$ such that $(x,y)\in\mathbb{R}^{k+m}$ and $\|(x,y)\|_{2}\leq 1.$ Then,
$\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\Big{\|}\geq\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\Big{\|}_{2}=\Big{\|}\begin{pmatrix}Ax\\ By\end{pmatrix}\Big{\|}_{2}=\sqrt{\|Ax\|_{2}^{2}+\|By\|_{2}^{2}}.$
The previous inequality is also true, if we choose $(x,y)=(\textbf{0},y)\in\mathbb{R}^{k+m}$ with $\|(\textbf{0},y)\|_{2}\leq 1.$ So,
$\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\Big{\|}\geq\|By\|_{2}.$
Taking the supremum for all $y\in\mathbb{R}^{m}$ with $\|y\|_{2}\leq 1,$ gives $\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\Big{\|}\geq\|B\|.$ Similarly, by choosing $(x,\textbf{0})\in\mathbb{R}^{k+m}$ with $\|(x,\textbf{0})\|_{2}\leq 1$ gives $\Big{\|}\begin{bmatrix}A&\textbf{0}_{1}\\ \textbf{0}_{2}&B\end{bmatrix}\Big{\|}\geq\|A\|,$ concluding the assertion. ∎
Consider the IFS $\mathcal{I}_{1}=\{\mathbb{R};h_{1},h_{2},\ldots,h_{N}\}$ and $\mathcal{I}_{2}=\{\mathbb{R};g_{1},g_{2},\ldots,g_{N}\}$ consists of similarity mappings on $\mathbb{R},$ i.e., $h_{i}(x)=c_{i}x+d_{i}$ and $g_{i}(x)=p_{i}x+q_{i},$ where $c_{i},p_{i}$ are similarity ratios in (1,1)(-1,1) and $d_{i},q_{i}\in\mathbb{R}$ are translation vectors. The product of IFS $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ is given as $\mathcal{I}_{1}\times\mathcal{I}_{2}$ and is defined as
$\mathcal{I}_{1}\times\mathcal{I}_{2}:=\{\mathbb{R}^{2};\ \psi_{ij}:i,j=1,2,\ldots,N\},$
where $\psi_{ij}(x,y):=(h_{i}(x),g_{j}(y)).$ Clearly, for $\psi_{ij}$ to be simlitudes, $c_{i}=d_{j}=c,\ \forall\ i,j.$ Consider an example $\mathcal{I}_{1}=\Big{\{}[0,1]:h_{1}(x)=\frac{x}{3},h_{2}(x)=\frac{x}{3}+\frac{1}{3}\Big{\}}$ and $\mathcal{I}_{2}=\Big{\{}[0,1]:g_{1}(x)=\frac{x}{2},g_{2}(x)=\frac{x}{2}+\frac{1}{2}\Big{\}}.$ Then, $\psi_{11}(x,y)=\Big{(}\frac{x}{3},\frac{y}{2}\Big{)}.$ Then,
$\|\psi_{11}(x,y)-\psi_{11}(x_{*},y_{*})\|_{2}=\Big{\|}\Big{(}\frac{x}{3},\frac{y}{2}\Big{)}-\Big{(}\frac{x_{*}}{3},\frac{y_{*}}{2}\Big{)}\Big{\|}_{2}\leq\frac{1}{2}\|(x,y)-(x_{*},y_{*})\|_{2}.$
Clearly, $\psi_{11}$ is a contraction on $\mathbb{R}^{2},$ but not a similitude.
Theorem 3.2.
Let $\mathcal{I}_{1}=\{\mathbb{R}^{k};h_{1},h_{2},\ldots,h_{N}\}$ and $\mathcal{I}_{2}=\{\mathbb{R}^{m};g_{1},g_{2},\ldots,g_{N}\}$ be two IFSs such that $h_{i}(x)=cU_{i}x+a_{i}$ and $g_{i}(x)=cV_{i}x+b_{i},$ where $c\in(-1,1),a_{i}\in\mathbb{R}^{k},b_{i}\in\mathbb{R}^{m},U_{i}$ and $V_{i}$ are $k\times k$ and $m\times m$ orthogonal matrices, respectively. If $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, then the product IFS $\mathcal{I}_{1}\times\mathcal{I}_{2}$ given as
$\mathcal{I}_{1}\times\mathcal{I}_{2}=\{\mathbb{R}^{k+m};\psi_{ij}:i,j=1,2,\ldots,N\},\ \psi_{ij}=(h_{i},g_{j})$
satisfies ESC.
Proof.
Since IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, there exist constants $\delta_{1}$ and $\delta_{2}>0,$ respectively such that
$\Delta_{n}^{\mathcal{I}_{1}}>\delta_{1}^{n}~~\text{ and }~~\Delta_{n}^{\mathcal{I}_{2}}>\delta_{2}^{n},$
for all $n\in\mathbb{N}.$ Notice that for $x\in\mathbb{R}^{k}$ and $y\in\mathbb{R}^{m},$ one gets
$\psi_{ij}(x,y)=(h_{i}(x),g_{j}(y))=(cU_{i}x+a_{i},cV_{j}y+b_{j})=c\begin{bmatrix}U_{i}&\textbf{0}_{1}\\ \textbf{0}_{2}&V_{j}\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\begin{pmatrix}a_{i}\\ b_{j}\end{pmatrix}$,
where $\textbf{0}_{1}$ and $\textbf{0}_{2}$ are zero matrices of order $k\times m$ and $m\times k,$ respectively. In view of Lemma 3.1, one gets
$d(\psi_{ij},\psi_{i^{\prime}j^{\prime}})$ $=\Big{\|}\begin{bmatrix}U_{i}&\textbf{0}_{1}\\ \textbf{0}_{2}&V_{j}\end{bmatrix}-\begin{bmatrix}U_{i^{\prime}}&\textbf{0}_{1}\\ \textbf{0}_{2}&V_{j^{\prime}}\end{bmatrix}\Big{\|}+\Big{\|}\begin{pmatrix}a_{i}\\ b_{j}\end{pmatrix}-\begin{pmatrix}a_{i^{\prime}}\\ b_{j^{\prime}}\end{pmatrix}\Big{\|}_{2}$
$=\Big{\|}\begin{bmatrix}U_{i}-U_{i^{\prime}}&\textbf{0}_{1}\\ \textbf{0}_{2}&V_{j}-V_{j^{\prime}}\end{bmatrix}\Big{\|}+\Big{\|}\begin{pmatrix}a_{i}-a_{i^{\prime}}\\ b_{j}-b_{j^{\prime}}\end{pmatrix}\Big{\|}_{2}$
$\geq\max\{\|U_{i}-U_{i^{\prime}}\|,\|V_{j}-V_{j^{\prime}}\|\}+\sqrt{\|a_{i}-a_{i^{\prime}}\|_{2}^{2}+\|b_{j}-b_{j^{\prime}}\|_{2}^{2}}$
$\geq\max\{\|U_{i}-U_{i^{\prime}}\|+\|a_{i}-a_{i^{\prime}}\|_{2},\|V_{j}-V_{j^{\prime}}\|+\|b_{j}-b_{j^{\prime}}\|_{2}\}$
$=\max\{d(h_{i},h_{i^{\prime}}),d(g_{j},g_{j^{\prime}})\}$
$\geq\max\{\Delta_{n}^{\mathcal{I}_{1}},\Delta_{n}^{\mathcal{I}_{2}}\}.$
Thus,
$\Delta_{n}$ $=\inf\{d(\psi_{ij},\psi_{i^{\prime}j^{\prime}}):ij,i^{\prime}j^{\prime}\in\{1,2,\ldots,N\}^{n},ij\neq i^{\prime}j^{\prime}\}$
$\geq\max\{\Delta_{n}^{\mathcal{I}_{1}},\Delta_{n}^{\mathcal{I}_{2}}\}>\min\{\delta_{1}^{n},\delta_{2}^{n}\}=\delta_{3}^{n},$
where $\delta_{3}=\min\{\delta_{1},\delta_{2}\},$ for all $n\in\mathbb{N},$ assuring that $\mathcal{I}_{1}\times\mathcal{I}_{2}$ satisfies ESC. ∎
Remark 3.3.
The previous theorem plays a considerable role in the context of extending the Hochman [18, Theorem 1.4] groundbreaking result to the product of IFSs. Meanwhile, Bárány and Verma [5] introduced a weak ESC; it remains an open question whether the result holds in that case as well.
Consider an IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ on $\mathbb{R}^{m}$ that consists of a finite number of affine mappings. Then, there exists a unique nonempty compact set $A\subset\mathbb{R}^{m}$ (termed as an invariant set) such that $A=\bigcup_{i=1}^{N}h_{i}(A).$ Let $\text{co}(A)$ be the closed convex hull of $A.$ Define $\Delta_{n}^{*}$ as
(3.1) $\Delta_{n}^{*}:=\inf\{\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A))):i,j\in\{1,2,\ldots,N\}^{n},i\neq j\}.$
Then, IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ is said to satisfy the modified ESC if there exists a $\delta>0$ such that $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$
Note 3.4.
It is noted that several researchers attempted to generalise the concept of ESC introduced by Hochman [17, 18]. In a way, Bárány and Verma [5] defined weak ESC for self-similar IFS, ESC for graph-directed self-similar IFS, ESC for common fixed point system. Above all that, the most fundamental advantage of the modified ESC is that it is defined not only for self-similar IFS but also extends to more general classes of IFS, including self-affine IFS and self-conformal IFS.
Example 3.5.
Some basic examples are presented for the reader’s convenience. Consider the middle third Cantor set $\mathcal{C}$ on $[0,1].$ Then, its IFS is given as $\{[0,1]:h_{1}(x)=\frac{x}{3},h_{2}(x)=\frac{x}{3}+\frac{2}{3}\}.$ Clearly,
$\mathcal{C}=h_{1}(\mathcal{C})\cup h_{2}(\mathcal{C}).$
Since the contraction ratios are the same for both maps in the IFS, only the translation vectors will play a role. For $i=11\ldots 1\in\{1,2\}^{n},$ we get $h_{i}(x)=\frac{x}{3^{n}},$ and for $i=22\ldots 2\in\{1,2\}^{n},$ we have $h_{i}(x)=\frac{1}{3^{n}}(x+2)+\frac{2}{3^{n-1}}+\cdots+\frac{2}{3}.$ Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2\}^{n}.$ Then, $h_{j}(x)=\frac{x}{3^{n}}+\frac{2}{3^{n}}$ and
$d(h_{i},h_{j})=\frac{2}{3^{n}}.$
It follows that
$\Delta_{n}=\inf\{d(h_{i},h_{j}):i,j\in\{1,2\}^{n},i\neq j\}=\frac{2}{3^{n}}>\frac{1}{3^{n}}.$
Clearly, $\delta=\frac{1}{3}.$ Further, the closed convex hull of the middle third Cantor set $\mathcal{C}$ is $[0,1].$ Then,
$\Delta_{n}^{*}$ $=\inf\{\mathfrak{H}(h_{i}(\text{co}(\mathcal{C})),h_{j}(\text{co}(\mathcal{C}))):i,j\in\{1,2\}^{n},i\neq j\}$
$=\inf\{\mathfrak{H}(h_{i}([0,1]),h_{j}([0,1])):i,j\in\{1,2\}^{n},i\neq j\}.$
Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2\}^{n}.$ Then, $h_{i}([0,1])=\big{[}0,\frac{1}{3^{n}}\big{]}$ and $h_{j}([0,1])=\big{[}\frac{2}{3^{n}},\frac{1}{3^{n-1}}\big{]}.$ Also
$\mathfrak{H}(h_{i}([0,1]),h_{j}([0,1]))=\mathfrak{H}\Big{(}\Big{[}0,\frac{1}{3^{n}}\Big{]},\Big{[}\frac{2}{3^{n}},\frac{1}{3^{n-1}}\Big{]}\Big{)}=\frac{2}{3^{n}},$
and
$\Delta_{n}^{*}=\inf\{\mathfrak{H}(h_{i}([0,1]),h_{j}([0,1])):i,j\in\{1,2\}^{n},i\neq j\}=\frac{2}{3^{n}}>\frac{1}{3^{n}}.$
The second example can be considered as Sierpiński gasket on $\mathbb{R}^{2}.$ Its IFS is elaborated as $\{\mathbb{R}^{2}:~h_{1},h_{2},h_{3}\},$ where
$h_{1}\begin{pmatrix}x\\ y\end{pmatrix}$ $=\frac{1}{2}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}$,
$h_{2}\begin{pmatrix}x\\ y\end{pmatrix}$ $=\frac{1}{2}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\begin{pmatrix}\frac{1}{2}\\ 0\end{pmatrix}$,
$h_{3}\begin{pmatrix}x\\ y\end{pmatrix}$ $=\frac{1}{2}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\begin{pmatrix}\frac{1}{4}\\ \frac{\sqrt{3}}{4}\end{pmatrix}$.
Now we will try to estimate
$\Delta_{n}=\inf\{d(h_{i},h_{j}):i,j\in\{1,2,3\}^{n},i\neq j\}.$
Choose $i,j\in\{1,2,3\}^{n},i\neq j$ as $i=11\ldots 1$ and $j=11\ldots 12.$ Then,
$h_{i}\begin{pmatrix}x\\ y\end{pmatrix}=\frac{1}{2^{n}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}~\text{ and }~h_{j}\begin{pmatrix}x\\ y\end{pmatrix}=\frac{1}{2^{n}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\begin{pmatrix}\frac{1}{2^{n}}\\ 0\end{pmatrix}$.
Thus, $d(h_{i},h_{j})=\frac{1}{2^{n}}$ and $\Delta_{n}=\frac{1}{2^{n}}.$ Now the closed convex hull of the Sierpiński gasket is an equilateral triangle $\triangle$ with vertices $(0,0),(1,0,),(\frac{1}{2},\frac{\sqrt{3}}{2}),$ so
$\Delta_{n}^{*}$ $=\inf\{\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle)):i,j\in\{1,2,3\}^{n},i\neq j\}.$
Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2,3\}^{n}.$ Then, $h_{i}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(0,0),(\frac{1}{2^{n}},0),(\frac{1}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}})$ and $h_{j}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(\frac{1}{2^{n}},0),(\frac{1}{2^{n-1}},0),\\ (\frac{3}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}}).$ Then, $\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle))=\frac{1}{2^{n}}$ and $\Delta_{n}^{*}=\frac{1}{2^{n}}.$
Remark 3.6.
It is remarkable that $\Delta_{n}=0,$ for all $n\in\mathbb{N}$ if and only if $\Delta_{n}^{*}=0,$ for all $n\in\mathbb{N}.$ Let $\Delta_{n}^{*}=0,$ for all $n\in\mathbb{N}.$ Since the infimum is taken over a finite set, there exist $i,j\in\{1,2,\ldots,N\}^{n}$ with $i\neq j$ such that
$\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))=0.$
It follows that $h_{i}=h_{j}$ on $co(A),$ that is, the maps $h_{i}$ and $h_{j}$ are same on the closed convex hull of the attractor. So, $d(h_{i},h_{j})=0,$ claiming that $\Delta_{n}=0,$ for all $n\in\mathbb{N}.$ On the same lines converse part is done.
Theorem 3.7.
Consider an IFS consisting of similarity transformations $(h_{i}(x)=c_{i}U_{i}x+a_{i},i=1,2,\ldots,N)$ on $\mathbb{R}^{m},$ then there exists a constant $k\in\mathbb{R}$ such that $\Delta_{n}^{*}\leq k\Delta_{n}.$
Proof.
Let $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots,N\}^{n}.$ Then,
$h_{i}(x)$ $=h_{i_{1}}\circ h_{i_{2}}\circ\dots\circ h_{i_{n}}(x)$
$=h_{i_{1}}\circ h_{i_{2}}\circ\dots\circ h_{i_{n-1}}(c_{i_{n}}U_{i_{n}}x+a_{i_{n}})$
$=h_{i_{1}}\circ h_{i_{2}}\circ\dots\circ h_{i_{n-2}}(c_{i_{n-1}}c_{i_{n}}U_{i_{n-1}}U_{i_{n}}x+c_{i_{n-1}}U_{i_{n-1}}a_{i_{n}}+a_{i_{n-1}})$
$=(c_{i_{1}}\dots c_{i_{n-1}}c_{i_{n}})(U_{i_{1}}\dots U_{i_{n}}x)+(c_{i_{1}}\ldots c_{i_{n-1}}U_{i_{1}}\ldots U_{i_{n-1}}a_{i_{n}}+$
$\quad c_{i_{1}}\ldots c_{i_{n-2}}U_{i_{1}}\ldots U_{i_{n-2}}a_{i_{n-1}}+\cdots+c_{i_{1}}U_{i_{1}}a_{i_{2}}+a_{i_{1}}).$
Let $c_{i}=c_{i_{1}}\dots c_{i_{n}},~U_{i}=U_{i_{1}}\dots U_{i_{n}},$ and $d_{i}=c_{i_{1}}\ldots c_{i_{n-1}}U_{i_{1}}\ldots U_{i_{n-1}}a_{i_{n}}+c_{i_{1}}\ldots c_{i_{n-2}}U_{i_{1}}\ldots U_{i_{n-2}}a_{i_{n-1}}\\ +\cdots+c_{i_{1}}U_{i_{1}}a_{i_{2}}+a_{i_{1}}.$ It follows that $h_{i}(x)=c_{i}U_{i}x+d_{i}.$ For $i,j\in\{1,2,\ldots,N\}^{n},$ we have
$d(h_{i},h_{j})=\left|\log c_{i}-\log c_{j}\right|+\left\|U_{i}-U_{j}\right\|+\left\|d_{i}-d_{j}\right\|_{2},~~\text{ and}$
$\mathfrak{H}(h_{i}(\text{co}(A)),$ $h_{j}(\text{co}(A)))=\max\Big{\{}\sup_{x\in\text{co}(A)}\inf_{y\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2},\sup_{y\in\text{co}(A)}\inf_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\Big{\}}.$
Now,
$\inf_{y\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\leq\|h_{i}(x)-h_{j}(x)\|_{2}.$
Taking supremum both side
$\sup_{x\in\text{co}(A)}\inf_{y\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\leq\sup_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(x)\|_{2}.$
Similarly,
$\sup_{y\in\text{co}(A)}\inf_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\leq\sup_{y\in\text{co}(A)}\|h_{i}(y)-h_{j}(y)\|_{2}.$
This gives
(3.2) $\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))\leq\sup_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(x)\|_{2}.$
Consider
$\|h_{i}(x)-h_{j}(x)\|_{2}$ $=\|c_{i}U_{i}x+d_{i}-(c_{j}U_{j}x+d_{j})\|_{2}$
$\leq\|(c_{i}U_{i}-c_{j}U_{j})x\|_{2}+\|d_{i}-d_{j}\|_{2}$
$\leq\|c_{i}U_{i}-c_{j}U_{j}\|\cdot\|x\|_{2}+\|d_{i}-d_{j}\|_{2}$
$\leq\|c_{i}U_{i}-c_{j}U_{j}\|\cdot k+\|d_{i}-d_{j}\|_{2},$
Where $k=\max\{\|x\|_{2}:x\in\text{co}(A)\}.$ Let $k_{*}=\max\{k,1\}.$
$\|h_{i}(x)-h_{j}(x)\|_{2}$ $\leq k_{*}(\|c_{i}U_{i}-c_{j}U_{j}\|+\|d_{i}-d_{j}\|_{2})$
$\leq k_{*}(|c_{i}-c_{j}|\|U_{i}\|+|c_{j}|\|U_{i}-U_{j}\|+\|d_{i}-d_{j}\|_{2})$
$\leq k_{*}(|c_{i}-c_{j}|+\|U_{i}-U_{j}\|+\|d_{i}-d_{j}\|_{2}),$
as $c_{j}<1,\ \forall~j\in\{1,2,\ldots,N\}^{n}.$ Using mean value theorem on interval $c_{i},c_{j}]$ if $c_{i}<c_{j},$ (otherwise $c_{j},c_{i}],$ we get $|c_{i}-c_{j}|<|\log c_{i}-\log c_{j}|.$ So,
$\|h_{i}(x)-h_{j}(x)\|_{2}$ $\leq k_{*}(|\log c_{i}-\log c_{j}|+\|U_{i}-U_{j}\|+\|d_{i}-d_{j}\|_{2})$
$\leq k_{*}\ d\left(h_{i},h_{j}\right).$
Consequently, $\sup_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(x)\|_{2}\leq k_{*}\ d(h_{i},h_{j}).$ Using equation (3.2), we get
$\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))\leq k_{*}\ d(h_{i},h_{j}),~~\forall~i,j.$
Since the constant $k_{*}$ is independent of $n$, taking infimum on both sides, we have
$\Delta_{n}^{*}\leq k_{*}\Delta_{n}.$
This completes the proof.
Proposition 3.8.
Consider a homogeneous IFS consisting of similarity transformations $(h_{i}(x)=cx+d_{i},i=1,2,\ldots,N)$ on $\mathbb{R},$ then the above two definitions (mentioned in equations (2.1) and (3.1)) of ESC coincide, that is, $\Delta_{n}=\Delta_{n}^{*}.$
Proof.
Consider $i,j\in\{1,2,\ldots,N\}^{n},$ we get
$d(h_{i},h_{j})=|d_{i}-d_{j}|.$
In $\mathbb{R},$ the closed convex hull of attractor $A$ is a compact interval $I=[a,b]$ say, then
$\mathfrak{H}(h_{i}(\text{co}(A),h_{j}(\text{co}(A)))$ $=\mathfrak{H}(h_{i}(I),h_{j}(I))$
$=\mathfrak{H}([ac+d_{i},bc+d_{i}],[ac+d_{j},bc+d_{j}])$
$=\max\{|ac+d_{i}-(ac+d_{j})|,|bc+d_{i}-(bc+d_{j})|\}$
$=|d_{i}-d_{j}|.$
Thus, in this case
$\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))=|d_{i}-d_{j}|=d(h_{i},h_{j}).$
Taking the infimum on both sides over $i,j\in\{1,2,\ldots,N\}^{n},$ we get $\Delta_{n}^{*}=\Delta_{n},$ concluding the proof.
Note 3.9.
If homogeneous IFS on $\mathbb{R}$ satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$ Thus, the homogeneous IFS satisfies modified ESC.
The next remark will give some hints on the ESC for the distinct contraction ratios.
Remark 3.10.
Consider an IFS on $\mathbb{R}$ with similarity mappings $(h_{i}(x)=c_{i}x+d_{i},~i=1,2,\ldots,N)$ with attractor $A.$ The closed convex hull of $A$ is a compact interval $I=[a,b],$ say. Consider the case $0\leq a<b,~c_{i}-c_{j}>0$ and $d_{i}-d_{j}\geq 0,$ for all $i,j\in\{1,2,\ldots,N\}^{n}.$ Suppose $i,j\in\{1,2,\ldots,N\}^{n},$ we get
$d(h_{i},h_{j})=|\log c_{i}-\log c_{j}|+|d_{i}-d_{j}|.$
The $\log$ function will satisfy
$\frac{|\log c_{i}-\log c_{j}|}{|c_{i}-c_{j}|}\leq\frac{1}{\min_{i\in\{1,2,\ldots,N\}^{n}}|c_{i}|}=\frac{1}{(\min_{i\in\{1,2,\ldots,N\}}|c_{i}|)^{n}}.$
Let $k_{n}=\frac{1}{(\min_{i\in\{1,2,\ldots,N\}}|c_{i}|)^{n}}.$ So,
$d(h_{i},h_{j})$ $\leq k_{n}|c_{i}-c_{j}|+|d_{i}-d_{j}|$
$\leq k_{n}\{(c_{i}-c_{j})+(d_{i}-d_{j})\}.$
Again, notice that
$\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))$ $=\mathfrak{H}(h_{i}(I),h_{j}(I))$
$=\mathfrak{H}([ac_{i}+d_{i},bc_{i}+d_{i}],[ac_{j}+d_{j},bc_{j}+d_{j}])$
$=\max\{|ac_{i}+d_{i}-(ac_{j}+d_{j})|,|bc_{i}+d_{i}-(bc_{j}+d_{j})|\}$
$=\max\{|a(c_{i}-c_{j})+(d_{i}-d_{j})|,|b(c_{i}-c_{j})+(d_{i}-d_{j})|\}.$
Without loss of generality, assume that
$\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))$ $=b(c_{i}-c_{j})+(d_{i}-d_{j})$
$\geq\min\{b,1\}\{(c_{i}-c_{j})+(d_{i}-d_{j})\},$
that is,
$\frac{1}{\min\{b,1\}}\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))\geq\{(c_{i}-c_{j})+(d_{i}-d_{j})\}\geq\frac{1}{k_{n}}d(h_{i},h_{j}).$
Taking infimum over $i,j\in\{1,2,\ldots,N\}^{n}$ both sides, we get $\frac{1}{\min\{b,1\}}\Delta_{n}^{*}\geq\frac{1}{k_{n}}\Delta_{n}.$ If the IFS satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\frac{\min\{b,1\}}{k_{n}}\delta^{n}=\min\{b,1\}\Big{(}\min_{i\in\{1,2,\ldots,N\}}|c_{i}|\delta\Big{)}^{n},$ for all $n\in\mathbb{N}.$
In the upcoming theorems, several dense subsets of $\mathcal{K}(\mathbb{R}^{m})$ are explored, which will be helpful in the dimension-preserving approximation of sets.
Theorem 3.11.
For each $\alpha\in[0,m],$ the set $A_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{A}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
Proof.
Let $\epsilon>0$ and $E\in\mathcal{K}(\mathbb{R}^{m}).$ Since $E$ is compact subset of $\mathbb{R}^{m},$ there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\bigcup_{j=1}^{n}B(e_{j},\epsilon),$ where $B(e_{j},\epsilon)$ is the open set centered at $e_{j}$ and radius $\epsilon.$ Subsequently, consider $F=\{e_{1},e_{2},\dots,e_{n}\}\in\mathcal{K}(\mathbb{R}^{m}).$ In view of definition of Hausdorff metric and Example 2.3, we get
$\mathfrak{H}(E,F)<\epsilon~~\text{ and }~~\dim_{\mathcal{A}}F=0.$
This shows that the assertion is true in the case of $\alpha=0.$ Again, assume $\alpha\in(0,m].$ Since the Assouad dimension satisfies the open set property (see Theorem 2.4), $\dim_{\mathcal{A}}B(e_{j},\epsilon)=m.$ Now using [8, Theorem 1], there exists a set $F_{e_{1}}\subset B(e_{1},\epsilon)$ such that $\dim_{\mathcal{A}}F_{e_{1}}=\alpha.$ Define $F^{\prime}=F\cup F_{e_{1}}.$ In view of the finite stability of the Assouad dimension, we get
$\dim_{\mathcal{A}}F^{\prime}=\max\{\dim_{\mathcal{A}}F,\dim_{\mathcal{A}}F_{e_{1}}\}=\alpha,$
and $\mathfrak{H}(E,F^{\prime})<\epsilon,$ conclude the claim. ∎
Theorem 3.12.
For each $\alpha\in[0,m],$ the set $H_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
Proof.
Proof follows on similar arguments as in the previous theorem, hence we avoid the details. ∎
Note that for any $E\subset\mathbb{R}^{m},$ we have $\dim_{\mathcal{H}}E\leq\dim_{\mathcal{A}}E.$ Thus, our next result is worth mentioning.
Theorem 3.13.
The set $\mathcal{D}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E\neq\dim_{\mathcal{A}}E\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
Proof.
In case of $m=1,$ using [14, Theorem 2.1.1], one can clearly show that $\mathcal{D}$ is a non-empty subset of $\mathcal{K}(\mathbb{R}^{m}).$ The Bedford-McMullen carpet (denoted as $\mathcal{B}\subset\mathbb{R}^{2}$) has the Hausdorff dimension strictly less than the Assouad dimension (see [15, Theorem 2.1]) is an example in case of $m=2.$ For a higher dimension $(m>2)$, consider the set $\mathcal{B}\times\{(\underset{m-2\text{ times}}{\underbrace{0,0,\dots,0}})\}=\{(x,y,0,0,\ldots,0)\in\mathbb{R}^{m}:(x,y)\in\mathcal{B}\},$ concluding the set $\mathcal{D}$ is non-empty. Further, assume $E\in\mathcal{K}(\mathbb{R}^{m})$ and $\epsilon>0.$ Then, there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\cup_{j=1}^{n}B(e_{j},\epsilon).$ Observe that the Hausdorff and the Assouad dimensions are invariant under scaling and translation (see [11, 14]). Using translation and scaling, we get $\mathcal{B}\times\{(0,0,\ldots,0)\}\subset B(e_{1},\epsilon).$ Choose $F=\{e_{1},e_{2},\ldots,e_{n}\}\cup(\mathcal{B}\times\{(0,0,\ldots,0)\}).$ Since the Assouad and the Hausdorff dimensions satisfy finite stability, we get
$\dim_{\mathcal{A}}F=\dim_{\mathcal{A}}(\mathcal{B}\times\{(0,0,\ldots,0)\}),$
$\dim_{\mathcal{H}}F=\dim_{\mathcal{H}}(\mathcal{B}\times\{(0,0,\ldots,0)\}).$
Clearly, $\dim_{\mathcal{A}}F\neq\dim_{\mathcal{H}}F$ and $\mathfrak{H}(E,F)<\epsilon,$ concluding the assertion. ∎
Before mentioning our upcoming theorem, we define the following for an IFS $\mathcal{I}=\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ with $h_{i}(x)=c_{i}U_{i}x+a_{i},$ where $0<c_{i}<1$ and $U_{i}$ is $m\times m$ orthogonal matrix, $a_{i}\in\mathbb{R}^{m}.$
  • An IFS $\mathcal{I}$ is said to be affinely irreducible if the only trivial affine subspaces are simultaneously preserved by all $h_{i}.$
  • A linear subspace $V\subseteq\mathbb{R}^{m}$ is said to be $D\mathcal{I}$-invariant if it is invariant under the orthogonal parts, i.e., $U_{i}(V)\subseteq V,$ for all $i.$
We now mention the following theorem, which is a conclusion of [18, Theorem 1.4].
Theorem 3.14.
Let $A\subset\mathbb{R}^{m}$ be the attractor of an affinely irreducible IFS given as $\mathcal{I}=\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ with $h_{i}(x)=c_{i}U_{i}x+a_{i},$ where $0<c_{i}<1$ and $U_{i}$ is $m\times m$ orthogonal matrix, $a_{i}\in\mathbb{R}^{m}.$ Suppose that $\mathcal{I}$ satisfies ESC and there does not exist a non-trivial $D\mathcal{I}$-invariant linear subspace $V\subseteq\mathbb{R}^{m}$ and $e\in A$ with $\dim_{\mathcal{H}}(A\cap(V+e))=\dim_{\mathcal{H}}V.$ Then, $\dim_{\mathcal{H}}A=\min\{m,\dim_{S}A\},$ where $\dim_{S}A$ is the similarity dimension of $A.$
Let $\Lambda=\bigcup_{n\in\mathbb{N}}\{1,2,\ldots,N\}^{n}$ be the collection of all finite alphabets on $\{1,2,\ldots,N\}.$
Theorem 3.15.
Let $\mathcal{I}$ be an IFS consisting of a finite number of similarity transformations (say, $h_{i}(x)=c_{i}U_{i}x+a_{i},~i=1,2,\ldots,N$) in $\mathbb{R}^{m}$ with attractor as $A.$ If there exists an alphabet $i\in\Lambda$ such that the IFS $\mathcal{I}_{n}=\{\mathbb{R}^{m};~h_{ji}:j\in\{1,2,\ldots,N\}^{n}\}$ is affinely irreducible with attractor $A_{n}$ satisfies ESC and there does not exist a non-trivial $D\mathcal{I}_{n}$-invariant linear subspace $V\subseteq\mathbb{R}^{m}$ and $e\in A_{n}$ with $\dim_{\mathcal{H}}(A_{n}\cap(V+e))=\dim_{\mathcal{H}}V.$ Then, $\dim_{\mathcal{H}}A=\min\{m,\dim_{S}A\},$ where $\dim_{S}A$ is the similarity dimension of $A.$
Proof.
The IFS $\mathcal{I}_{n}$ satisfies the assumptions of Theorem 3.14, so $\dim_{\mathcal{H}}A_{n}=\min\{m,\dim_{S}A_{n}\},$ where $\sum_{j\in\{1,2,\ldots,N\}^{n}}c_{ji}^{\dim_{S}A_{n}}=1.$ Using the fact $A_{n}\subseteq A,$ we get $\dim_{\mathcal{H}}A_{n}\leq\dim_{\mathcal{H}}A.$ In the first case $\dim_{\mathcal{H}}A_{n}=m,$ then $\dim_{\mathcal{H}}A=m,$ we are done. In another case, consider $\dim_{\mathcal{H}}A_{n}=\dim_{S}A_{n}.$ On the contrary, suppose that $\dim_{\mathcal{H}}A<\dim_{S}A.$ Then,
$c_{i}^{-\dim_{S}A_{n}}=\sum_{j\in\{1,2,\ldots,N\}^{n}}c_{j}^{\dim_{S}A_{n}}\geq\sum_{j\in\{1,2,\ldots,N\}^{n}}c_{j}^{\dim_{\mathcal{H}}A}=$
$\sum_{j\in\{1,2,\ldots,N\}^{n}}c_{j}^{(\dim_{\mathcal{H}}A-\dim_{S}A)}c_{j}^{\dim_{S}A}\geq\sum_{j\in\{1,2,\ldots,N\}^{n}}c_{\max}^{n(\dim_{\mathcal{H}}A-\dim_{S}A)}c_{j}^{\dim_{S}A}$
$=c_{\max}^{n(\dim_{\mathcal{H}}A-\dim_{S}A)}.$
As $n\rightarrow\infty,$ we get a contradiction. Thus, $\dim_{\mathcal{H}}A=\dim_{S}A,$ concluding the proof. ∎
Remark 3.16.
It is straightforward to observe that if an IFS satisfies ESC (SSC or OSC or SOSC) at the first level, then all subcollections of the IFS at $n$-th level also satisfy ESC (SSC or OSC or SOSC). But the converse is not true in general. That is, if there exists a subcollection of the IFS of the $n$-th level that satisfies ESC (SSC or OSC or SOSC), then the original IFS does not necessarily satisfy ESC (SSC or OSC or SOSC). As an example, consider an IFS on $\mathbb{R}$ with three mappings as follows:
$h_{1}(x)=\frac{x}{4},\quad h_{2}(x)=\frac{x}{4}+\frac{9}{16},\quad h_{3}(x)=\frac{x}{4}+\frac{3}{4}.$
It is noted that the IFS does not satisfy SSC, but there exists a subcollection of $2$-nd level IFS satisfying SSC (see [17, 35]). In view of this, the previous theorem is a weaker version of the Hochman result [18, Theorem 1.4].
To our knowledge, we could not find any proof or statement of the following lemma in the literature; however, some partial mentions for Dirac measures can be seen in [30] and also, since convolution is the smoothing operation, we believe $\dim_{L^{q}}(\mu*\nu)\geq\dim_{L^{q}}(\mu).$ So, it is crucial to include the complete proof with all details for the reader’s convenience.
Lemma 3.17.
Let $\mathcal{S}(\mathbb{R}^{m})=\big{\{}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}:x_{j}\in\mathbb{R}^{m},c_{j}>0,\sum_{j=1}^{n}c_{j}=1,n\in\mathbb{N}\big{\}}$ and $\mu$ be a probability measure with finite support. Then, for a fixed $\omega\in\mathcal{S}(\mathbb{R}^{m}),$
$\dim_{L^{q}}(\mu*\omega)=\dim_{L^{q}}(\mu),$
where $\mu*\omega$ is the convolution of $\mu$ and $\omega.$
Proof.
We first observe our claim for the Dirac measures. For this, consider $x\in\mathbb{R}$ with $x$ as dyadic end point, i.e., $x=\frac{j_{0}}{2^{k}}.$ Then,
$I-x=\Big{[}\frac{j}{2^{k}}-\frac{j_{0}}{2^{k}},\frac{j+1}{2^{k}}-\frac{j_{0}}{2^{k}}\Big{)}=\Big{[}\frac{j-j_{0}}{2^{k}},\frac{j+1-j_{0}}{2^{k}}\Big{)}.$
In this case, IxI-x is another subinterval in $D_{k}.$ So, $\dim_{L^{q}}(\mu*\delta_{x})=\dim_{L^{q}}(\mu).$ In second case observe that for any $x\in\mathbb{R},$ consider
$I-x=\Big{[}\frac{j}{2^{k}}-x,\frac{j+1}{2^{k}}-x\Big{)}.$
Note that the length of IxI-x is the same as the length of $I,$ equal to $\frac{1}{2^{k}}.$ So. IxI-x must be contained in two consecutive intervals (say, $I^{\prime}$ and $I^{\prime\prime}$) of $D_{k}$.
$\sum_{I\in D_{k}}\Big{(}\mu*\delta_{x}(I)\Big{)}^{q}=\sum_{I\in D_{k}}\Big{(}\mu(I-x)\Big{)}^{q}\leq\sum_{I^{\prime},I^{\prime\prime}\in D_{k}}\Big{(}\mu(I^{\prime})+\mu(I^{\prime\prime})\Big{)}^{q},$
where $I-x\subset I^{\prime}\cup I^{\prime\prime}.$
$\sum_{I\in D_{k}}\Big{(}\mu*\delta_{x}(I)\Big{)}^{q}\leq\sum_{I^{\prime},I^{\prime\prime}\in D_{k}}2^{q-1}\Big{(}\mu(I^{\prime})^{q}+\mu(I^{\prime\prime})^{q}\Big{)}\leq 2^{q}\sum_{I\in D_{k}}\mu(I)^{q}.$
Consequently,
$-\log\sum_{I\in D_{k}}\Big{(}\mu*\delta_{x}(I)\Big{)}^{q}\geq-\log\Big{(}2^{q}\sum_{I\in D_{k}}\mu(I)^{q}\Big{)}=-\log(2^{q})-\log\Big{(}\sum_{I\in D_{k}}\mu(I)^{q}\Big{)}.$
Dividing by $k$ and taking the limit infimum, we get
$\tau(\mu*\delta_{x},q)\geq\tau(\mu,q)~\text{ and }~\dim_{L^{q}}(\mu*\delta_{x})\geq\dim_{L^{q}}(\mu).$
Again, assume $I$ and $J$ are distinct intervals in $D_{k}.$ Then, for any $x\in\mathbb{R},~I-x$ and JxJ-x are disjoint because for $z\in(I-x)\cap(J-x)$, it follows z=yxz=y-x and $z=y^{\prime}-x$ for some $y\in I$ and $y^{\prime}\in J.$ So $x+z\in I\cap J,$ which is a contradiction. Thus, the collection $\{I-x:I\in D_{k}\}$ contains mutually disjoint intervals. For any $I\in D_{k},$ we have
$I\subset(I^{\prime}-x)\cup(I^{\prime\prime}-x)~~\text{ for some }I^{\prime},I^{\prime\prime}\in D_{k}.$
Following the same lines as in the previous one, we get $\dim_{L^{q}}(\mu*\delta_{x})\leq\dim_{L^{q}}(\mu).$ Let $\omega\in\mathcal{S}(\mathbb{R}^{m})$ be of the form $\omega=\sum_{j=1}^{n}c_{j}\delta_{x_{j}},$ for some $x_{j}\in\mathbb{R}^{m},c_{j}>0$ and $\sum_{j=1}^{n}c_{j}=1.$ Then, for $I\in D_{k},$ we have
$\Big{(}\Big{(}\mu*\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big{)}(I)\Big{)}^{q}$ $=\Big{(}\Big{(}\sum_{j=1}^{n}c_{j}\mu*\delta_{x_{j}}\Big{)}(I)\Big{)}^{q}$
$=\Big{(}\sum_{j=1}^{n}c_{j}^{q^{\prime}}\Big{)}^{\frac{q}{q^{\prime}}}\Big{(}\sum_{j=1}^{n}(\mu*\delta_{x_{j}}(I))^{q}\Big{)},$
where $q^{\prime}>1$ such that $\frac{1}{q^{\prime}}+\frac{1}{q}=1.$ Let $M=(\sum_{j=1}^{n}c_{j}^{q^{\prime}}\Big{)}^{\frac{q}{q^{\prime}}}.$ Thus,
$\sum_{I\in D_{k}}\Big{(}\Big{(}\mu*\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big{)}(I)\Big{)}^{q}$ $=M\sum_{I\in D_{k}}\Big{(}\sum_{j=1}^{n}(\mu*\delta_{x_{j}}(I))^{q}\Big{)}$
$=M\sum_{j=1}^{n}\Big{(}\sum_{I\in D_{k}}(\mu*\delta_{x_{j}}(I))^{q}\Big{)}$
$\leq 2^{q}Mn\Big{(}\sum_{I\in D_{k}}\mu(I)^{q}\Big{)}.$
This gives $\tau(\mu*\sum_{j=1}^{n}c_{j}\delta_{x_{j}},q)\geq\tau(\mu,q).$ Again,
$\sum_{I\in D_{k}}\mu(I)^{q}=\sum_{I\in D_{k}}\sum_{j=1}^{n}c_{j}\mu(I)^{q}=\sum_{I\in D_{k}}c_{1}\mu(I)^{q}+\cdots+\sum_{I\in D_{k}}c_{n}\mu(I)^{q}\leq$
$2^{q}\sum_{I\in D_{k}}c_{1}\Big{(}\mu*\delta_{x_{1}}(I)\Big{)}^{q}+\cdots+2^{q}\sum_{I\in D_{k}}c_{n}\Big{(}\mu*\delta_{x_{n}}(I)\Big{)}^{q}\leq 2^{q}\sum_{I\in D_{k}}\Big{(}c_{1}\mu*\delta_{x_{1}}(I)\Big{)}^{q}$
$+\cdots+2^{q}\sum_{I\in D_{k}}\Big{(}c_{n}\mu*\delta_{x_{n}}(I)\Big{)}^{q}\leq 2^{q}\sum_{I\in D_{k}}\Big{(}\sum_{j=1}^{n}c_{j}\mu*\delta_{x_{j}}(I)\Big{)}^{q}=2^{q}\sum_{I\in D_{k}}\Big{(}\mu*\sum_{j=1}^{n}c_{j}\delta_{x_{j}}(I)\Big{)}^{q},$
concluding that $\tau(\mu*\sum_{j=1}^{n}c_{j}\delta_{x_{j}},q)\leq\tau(\mu,q).$ This completes the proof. ∎
Lemma 3.18.
Let $\mu$ be a Borel probability measure on $\mathbb{R}^{m}$ and $\beta$ be a non-zero real number. Then, $\dim_{L^{q}}(\mu)=\dim_{L^{q}}(\nu),$ where $\nu(D)=\mu(\beta D),$ for all Borel set $D.$
Proof.
The result will be demonstrated in $\mathbb{R}$ and following a similar approach, we extend it to $\mathbb{R}^{m}.$ Firstly, suppose $\beta$ as a dyadic end point, that is, $\beta=\frac{j_{0}}{2^{s}},$ with $s>0.$ Then, $\beta I=\Big{[}\frac{jj_{0}}{2^{k+s}},\frac{jj_{0}+j_{0}}{2^{k+s}}\Big{)},$ and its length is equal to $\frac{j_{0}}{2^{k+s}}.$ Notice that
$\beta I=\Big{[}\frac{jj_{0}}{2^{k+s}},\frac{jj_{0}+1}{2^{k+s}}\Big{)}\cup\Big{[}\frac{jj_{0}+1}{2^{k+s}},\frac{jj_{0}+2}{2^{k+s}}\Big{)}\cup\cdots\cup\Big{[}\frac{jj_{0}+j_{0}-1}{2^{k+s}},\frac{jj_{0}+j_{0}}{2^{k+s}}\Big{)}.$
Consider $I_{1}=\Big{[}\frac{jj_{0}}{2^{k+s}},\frac{jj_{0}+1}{2^{k+s}}\Big{)},\ldots,I_{j_{0}}=\Big{[}\frac{jj_{0}+j_{0}-1}{2^{k+s}},\frac{jj_{0}+j_{0}}{2^{k+s}}\Big{)},$ so $\mu(\beta I)=\mu(I_{1})+\cdots+\mu(I_{j_{0}}).$ Thus,
$\sum_{I\in D_{k}}\nu(I)^{q}$ $=\sum_{I\in D_{k}}\mu(\beta I)^{q}=\sum_{I_{1},\ldots,I_{j_{0}}\in D_{k+s}}\Big{(}\mu(I_{1})+\cdots+\mu(I_{j_{0}})\Big{)}^{q}$
$\leq C\Big{\{}\sum_{I_{1}\in D_{k+s}}\mu(I_{1})^{q}+\cdots+\sum_{I_{j_{0}}\in D_{k+s}}\mu(I_{j_{0}})^{q}\Big{\}}=j_{0}C\sum_{I\in D_{k+s}}\mu(I)^{q},$
where $C$ is a suitable constant. It follows that
$\frac{-\log\sum_{I\in D_{k}}\nu(I)^{q}}{k}\geq\frac{-\log\Big{(}j_{0}2^{q-1}\sum_{I\in D_{k+s}}\mu(I)^{q}\Big{)}}{k+s}.$
Applying $\liminf_{k\rightarrow\infty}$ on both sides, we get $\dim_{L^{q}}(\nu)\geq\dim_{L^{q}}(\mu).$ For reverse inequality, we have
$I=\Big{[}\frac{j}{2^{k}},\frac{j+1}{2^{k}}\Big{)}=\Big{[}\frac{\beta j2^{s}}{j_{0}2^{k}},\frac{\beta(j+1)2^{s}}{j_{0}2^{k}}\Big{)}=\beta\Big{[}\frac{j}{j_{0}2^{k-s}},\frac{(j+1)}{j_{0}2^{k-s}}\Big{)}.$
Suppose that the interval $\Big{[}\frac{j}{j_{0}2^{k-s}},\frac{(j+1)}{j_{0}2^{k-s}}\Big{)}$ is contained in intervals $I_{1},I_{2},\ldots,I_{l}$ dyadic intervals of level ks.k-s. Then,
$I\subset\beta I_{1}\cup\beta I_{2}\cup\ldots\cup\beta I_{l},$
and $\mu(I)\leq\mu(\beta I_{1})+\mu(\beta I_{2})+\cdots+\mu(\beta I_{l})=\nu(I_{1})+\nu(I_{2})+\cdots+\nu(I_{l}).$ Subsequently,
$\sum_{I\in D_{k}}\mu(I)^{q}\leq\sum_{I_{1},I_{2},\ldots,I_{l}\in D_{k-s}}\Big{(}\nu(I_{1})+\nu(I_{2})+\cdots+\nu(I_{l})\Big{)}^{q}\leq C\sum_{I\in D_{k-s}}\nu(I)^{q}$
$\implies\hskip 28.45274pt\frac{-\log\sum_{I\in D_{k}}\mu(I)^{q}}{k}\geq\frac{-\log\Big{(}C\sum_{I\in D_{k-s}}\nu(I)^{q}\Big{)}}{k-s},$
where $C$ is a suitable constant. Applying $\liminf_{k\rightarrow\infty}$ on both sides, we get $\dim_{L^{q}}(\mu)\geq\dim_{L^{q}}(\nu).$ In another case, suppose that $\beta$ is not a dyadic endpoint. Then, for some $s>0,$ we have
$\frac{j_{0}}{2^{s}}<\beta<\frac{j_{0}+1}{2^{s}}.$
This gives
$\beta I=\beta\Big{[}\frac{j}{2^{k}},\frac{j+1}{2^{k}}\Big{)}\subset\Big{[}\frac{jj_{0}}{2^{k+s}},\frac{jj_{0}+j_{0}}{2^{k+s}}\Big{)}\cup\Big{[}\frac{jj_{0}+j_{0}}{2^{k+s}},\frac{jj_{0}+j}{2^{k+s}}\Big{)}\cup\Big{[}\frac{jj_{0}+j}{2^{k+s}},\frac{jj_{0}+j+j_{0}+1}{2^{k+s}}\Big{)},$
that is, $\beta I$ is contained in the finite number of dyadic intervals of level $k+s.$ Following the same lines, as done in the first case, we get $\dim_{L^{q}}(\nu)\geq\dim_{L^{q}}(\mu).$ Further,
$I=\Big{[}\frac{j}{2^{k}},\frac{j+1}{2^{k}}\Big{)}=\beta\Big{[}\frac{j}{\beta 2^{k}},\frac{j+1}{\beta 2^{k}}\Big{)}.$
Since $\frac{2^{s}}{j_{0}+1}<\frac{1}{\beta}<\frac{2^{s}}{j_{0}},$ we have
$\Big{[}\frac{j}{\beta 2^{k}},\frac{j+1}{\beta 2^{k}}\Big{)}\subset\Big{[}\frac{j}{(j_{0}+1)2^{k-s}},\frac{j+1}{(j_{0}+1)2^{k-s}}\Big{)}\cup\Big{[}\frac{j+1}{(j_{0}+1)2^{k-s}},\frac{j}{j_{0}2^{k-s}}\Big{)}\cup\Big{[}\frac{j}{j_{0}2^{k-s}},\frac{j+1}{j_{0}2^{k-s}}\Big{)}.$
This shows that the interval $\Big{[}\frac{j}{\beta 2^{k}},\frac{j+1}{\beta 2^{k}}\Big{)}$ is contained in some finite number of dyadic intervals of level ks,k-s, suppose $I_{1},I_{2},\ldots,I_{l_{*}}.$ Consequently,
$I\subset\beta I_{1}\cup\beta I_{2}\cup\ldots\cup\beta I_{l_{*}}.$
Following the similar approach, we get $\dim_{L^{q}}(\mu)\geq\dim_{L^{q}}(\nu).$ This completes the proof.
Theorem 3.19.
For each $\beta\in[0,m]$, the set $\mathcal{S}_{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$.
Proof.
Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ and $\epsilon>0.$ Using the density of $\mathcal{S}(\mathbb{R}^{m})$ in $\mathcal{P}(\mathbb{R}^{m})$ (see [7]), there exists $\nu=\sum_{j=1}^{n}c_{j}\delta_{x_{j}}$ in $\mathcal{S}(\mathbb{R}^{m})$ such that
$d_{L}(\mu,\nu)<\frac{\epsilon}{2}.$
Further, we consider a non-zero compactly supported measure $\lambda\in\mathcal{S}_{\beta}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Lemmas 3.17 and 3.18 implies that $\dim_{L^{q}}(\lambda^{\prime})=\dim_{L^{q}}(\lambda_{1})=\dim_{L^{q}}(\lambda)=\beta,$ that is, $\lambda^{\prime}\in\mathcal{S}_{\beta}.$ In view of the properties of convolution mentioned in Section 2, we get
$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\nu-\int_{\mathbb{R}^{m}}fd\lambda^{\prime}\Big{|}\Big{\}}$
$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\Big{(}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big{)}-\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})\Big{|}\Big{\}}.$
Notice that
$\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})$ $=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\nu(x)d\lambda_{1}(y)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d(\sum_{j=1}^{n}c_{j}\delta_{x_{j}})(x)d\lambda_{1}(y)$
$=\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j}+y)d\lambda_{1}(y).$
Thus,
$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}f(x_{j})-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}\hskip 11.38092pt$
$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j})d\lambda_{1}(y)-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}|f(x_{j})-f(y+x_{j})|d\lambda_{1}(y)\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}.$
We claim that the support of $\lambda_{1}$ is a subset of $\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ Let $C\subset\mathbb{R}^{m}\setminus\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ If $x\in C,$ then $\|x\|_{2}>\frac{\epsilon}{4\sqrt{m}}$ and $\frac{8\sqrt{m}diam(K)\|x\|_{2}}{\epsilon}>2diam(K).$ So, $\lambda_{1}(C)=\lambda(\frac{8\sqrt{m}diam(K)C}{\epsilon})=0.$ The previous estimation is reduced to
$d_{L}(\nu,\lambda^{\prime})$ $\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}$
$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\frac{\epsilon}{2}d\lambda_{1}(y)\Big{\}}\leq\frac{\epsilon}{2}.$
Using the above estimate, we immediately get
$d_{L}(\mu,\lambda^{\prime})\leq d_{L}(\mu,\nu)+d_{L}(\nu,\lambda^{\prime})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon,$
showing that $\mathcal{S}_{\beta}$ is dense.
Definition 3.20.
[22] The Fourier transform of $\mu\in\mathcal{P}(\mathbb{R}^{m})$ is defined as
$\hat{\mu}(\xi)=\int e^{-2\pi i\xi.x}d\mu(x),~~~~\xi\in\mathbb{R}^{m}.$
Recall that the Fourier transform of convolution $\mu*\nu$ has the property $\widehat{\mu*\nu}(\xi)=\hat{\mu}(\xi)\hat{\nu}(\xi),$ for details; see [22]. It is said that $\mu$ is a Rajchman measure if $\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$ The Riemann-Lebesgue lemma says that $\mu$ is Rajchman whenever it is absolutely continuous with respect to the Lebesgue measure; refer to [27], and the references cited therein. Solomyak [32, 33] studied the Fourier decay for self-similar measures and homogeneous self-affine measures.
Proposition 3.21.
If the Fourier transform of a measure $\nu$ is bounded and $\mu$ is Rajchman (with power Fourier decay), then $\mu*\nu$ is also a Rajchman (with power Fourier decay). For any Rajchman measure $\mu$ and real number $\beta>0$, we have $\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=0$, where $\nu(D)=\mu(\beta D)$ for each Borel set $D$.
Proof.
Since $\nu$ is bounded, $\nu(\mathbb{R}^{m})<\infty.$ Consider
$\lim_{|\xi|\to\infty}|\widehat{\mu*\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)||\hat{\nu}(\xi)|=K\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$
Again, note that
$\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.x}d\nu(x)\Big{|}=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.\beta x}d\mu(x)\Big{|}=0,$
as $\mu$ is Rajchman, concluding the assertion. ∎
Theorem 3.22.
The set $\mathcal{S}_{R}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$. Moreover, for each $\beta\in[0,m]$, the sets $\mathcal{S}_{R}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ and $\mathcal{S}_{F}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\hat{\mu}(\xi)\text{ has power Fourier decay}\}$ are dense in $\mathcal{P}(\mathbb{R}^{m})$.
Proof.
Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ and $\epsilon>0.$ Using the density of $\mathcal{S}(\mathbb{R}^{m})$ in $\mathcal{P}(\mathbb{R}^{m})$, there exists $\nu=\sum_{j=1}^{n}c_{j}\delta_{x_{j}}$ in $\mathcal{S}(\mathbb{R}^{m})$ such that
$d_{L}(\mu,\nu)<\frac{\epsilon}{2}.$
Choose a non-zero compactly supported Rajchman measure $\lambda\in\mathcal{S}_{R}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Proposition 3.21 implies that $\lambda^{\prime}\in\mathcal{S}_{R},$ that is, $\lambda^{\prime}$ is Rajchman. Following the lines of Theorem 3.19, one gets $d_{L}(\mu,\lambda^{\prime})<\epsilon,$ showing that $\mathcal{S}_{R}$ is dense in $\mathcal{P}(\mathbb{R}^{m}).$ Further, to show $\mathcal{S}_{R}^{\beta}$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ choose $\lambda\in\mathcal{S}_{R}^{\beta}.$ This concludes the proof. ∎
4. Conclusion & future remarks
[…]
Theorem 4.1.
Let $0\leq\alpha\leq m$ and $\mu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, there exist Borel probability measures $\nu,\lambda$ such that $\mu=\nu*\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$
Theorem 4.2.
Let $0\leq\alpha\leq m$, and $\mu$ be a finite Borel measure such that $\dim_{L^{q}}(\mu)\leq\alpha.$ Then, there exist finite Borel measures $\nu,\lambda$ such that $\mu=\nu+\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$
Statements and Declarations
The authors have no competing interests to declare that are relevant to the content of this article. There are no data associated with this paper.
Acknowledgements
The first author is supported by the SEED grant project of IIIT Allahabad. The second author is financially supported by the Ministry of Education at the IIIT Allahabad. Some parts of the paper have been presented in lectures delivered by the first author at the workshop “Course on Exponential separation and $L^{q}$ dimension” organized by IIIT Allahabad on April 23, 2025.

### v2

@@
 Abstract.
-In the dimension theory of sets and measures, a recent breakthrough happened due to Hochman, who introduced the exponential separation condition and proved the Hausdorff dimension result for invariant sets and measures generated by similarities on the real line.
+In the dimension theory of sets and measures, a recent breakthrough happened due to Hochman, who introduced the exponential separation condition (ESC) and proved the Hausdorff dimension result for invariant sets and measures generated by similarities on the real line.
 Following this groundbreaking work, we make a modest contribution by weakening the condition.
-We also define some sets in the class of all nonempty compact sets using the Assouad and Hausdorff dimensions and subsets of measures in the Borel probability measures using the $L^{q}$ dimension and prove their density in the respective spaces.
+Further, we define the modified ESC using the convex hull of the attractor and show that for homogeneous self-similar IFS on $\mathbb{R},$ both definitions coincide.
+We also define some sets in the class of all nonempty compact sets using the Assouad and Hausdorff dimensions and subsets of measures in the space of Borel probability measures on $\mathbb{R}^{m}$ using the $L^{q}$ dimension and the Rajchman property, and prove their density in the respective spaces.
 Key words and phrases:
@@
   • Under OSC, $\dim_{\mathcal{H}}A=\min\{m,\dim_{S}A\},$ see [21, 26].
-  • IFS with bi-Lipschitz mappings $(s_{i}\|x-y\|_{2}\leq\|f_{i}(x)-f_{i}(y)\|_{2}\leq t_{i}\|x-y\|_{2})$ on $\mathbb{R}^{m}$ and SSC, we have $s\leq\dim_{\mathcal{H}}A\leq t,$ where $s$ and $t$ satisfies $\sum_{i=1}^{N}s_{i}^{s}=1$ and $\sum_{i=1}^{N}t_{i}^{t}=1,$ respectively, see [11, Proposition 9.6-9.7].
+  • IFS with bi-Lipschitz mappings $(s_{i}\|x-y\|_{2}\leq\|f_{i}(x)-f_{i}(y)\|_{2}\leq t_{i}\|x-y\|_{2})$ on $\mathbb{R}^{m}$ and SSC, we have $s\leq\dim_{\mathcal{H}}A\leq t,$ where $s$ and $t$ satisfies $\sum_{i=1}^{N}s_{i}^{s}=1$ and $\sum_{i=1}^{N}t_{i}^{t}=1,$ respectively, see [12, Proposition 9.6-9.7].
 It is noted that the upper bound holds without any separation condition, but the lower bound does not hold even if we assume OSC.
 Additionally, it does not serve as a bound for the Assouad dimension.
-For example, visit [14, Section 8.5].
+For example, visit [15, Section 8.5].
   • Total overlap implies $\dim_{\mathcal{H}}A<\min\{m,\dim_{S}A\}.$
-  • Generally, $\dim_{\mathcal{H}}A\leq\dim_{\mathcal{A}}A,$ but if $A$ is Ahlfors regular, then $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A,$ see [14, Theorem 6.4.1].
+  • Generally, $\dim_{\mathcal{H}}A\leq\dim_{\mathcal{A}}A,$ but if $A$ is Ahlfors regular, then $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A,$ see [15, Theorem 6.4.1].
 IFS with similitudes on $\mathbb{R},$ either $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A$ or $\dim_{\mathcal{A}}A=1.$
 […]
-  • In $\mathbb{R}$ with similarities and ESC, Shmerkin computed $\dim_{L^{q}}\mu=\min\{1,\tilde{\tau}(q)/(q-1)\},$ see [30, Theorem 6.6].
-  • IFS with similarity and OSC implies $\dim_{\mathcal{H}}\mu=\min\{m,\dim_{S}\mu\}$, see [10]. In this case, under SSC, Fraser and Howroyd computed $\dim_{\mathcal{A}}\mu,$ see [14, Theorem 7.4.1].
-It is well known that the Hausdorff dimension decreases under Lipschitz mapping ([11, Corollary 2.4]), but this is not the case with the Assouad dimension ([14, Corollary 7.2.2]).
+  • In $\mathbb{R}$ with similarities and ESC, Shmerkin computed $\dim_{L^{q}}\mu=\min\{1,\tilde{\tau}(q)/(q-1)\},$ see [38, Theorem 6.6].
+  • IFS with similarity and OSC implies $\dim_{\mathcal{H}}\mu=\min\{m,\dim_{S}\mu\}$, see [11]. In this case, under SSC, Fraser and Howroyd computed $\dim_{\mathcal{A}}\mu,$ see [15, Theorem 7.4.1].
+It is well known that the Hausdorff dimension decreases under Lipschitz mapping ([12, Corollary 2.4]), but this is not the case with the Assouad dimension ([15, Corollary 7.2.2]).
 Bedford [7] in his PhD thesis, studied the box dimension of self-similar sets.
@@
 Several recent works in the direction of estimation of the Hausdorff dimension of self-affine sets and measures can be found in [5, 13, 20, 24, 39].
+In geometric measure theory, the Fourier transform has emerged as a powerful technique for determining the Hausdorff dimension.
+The interplay between dimension theory and Fourier analysis was first observed by Kaufman [23], who proved one part of Marstrand’s projection theorem using the Fourier transform method.
+Furthermore, this technique allows researchers to introduce the Fourier dimension $(\dim_{\mathcal{F}})$ of any subset of $\mathbb{R}^{m}.$ For any Borel subset $A$ of $\mathbb{R}^{m},$ we have $\dim_{\mathcal{F}}(A)\leq\dim_{\mathcal{H}}(A)$, and $\dim_{\mathcal{F}}(A)=\dim_{\mathcal{H}}(A)$ iff $A$ is Salem set.
+Interested readers may refer to the book [25], and the references therein for details.
+Recent work on the Fourier decay of self-similar measures and homogeneous self-affine measures can be referred to [40, 41].
 The primary theorems in this article include the following.
-  • The weaker version of the Hochman result [18, Theorem 1.4] is shown in Theorem 3.15.
+  • The weaker version of the Hochman result [19, Theorem 1.4] is shown in Theorem 3.14.
   • The modified ESC is being defined using the Hausdorff metric and compared with the existing ESC, along with some interesting remarks.
@@
 2. Preliminaries
-Let $\mathcal{K}(\mathbb{R}^{m})$ denote the collection of all nonempty compact subsets of $\mathbb{R}^{m}$ in the Euclidean norm $\|.\|_{2}.$ The Hausdorff distance $\mathfrak{H}$ between any $A,B\in\mathcal{K}(\mathbb{R}^{m})$ is elaborated as
+Let $\chi(\mathbb{R}^{m})$ denote the collection of all nonempty compact subsets of $\mathbb{R}^{m}$ in the Euclidean norm $\|.\|_{2}.$ The Hausdorff distance $\mathfrak{H}$ between any $A,B\in\chi(\mathbb{R}^{m})$ is elaborated as
 $\mathfrak{H}(A,B):=\inf\{\epsilon>0:A\subset B_{\epsilon}\text{ and }B\subset A_{\epsilon}\},$
@@
 $\mathfrak{H}(A,B):=\max\Big\{\sup_{a\in A}\inf_{b\in B}\|a-b\|_{2},\sup_{b\in B}\inf_{a\in A}\|a-b\|_{2}\Big\}.$
+Definition 2.1.
+[25] The Fourier transform of any Borel probability measure $\eta$ in $\mathbb{R}^{m}$ is defined as
+$\hat{\eta}(\zeta)=\int e^{-2\pi i\zeta.y}d\eta(y),~~~~\zeta\in\mathbb{R}^{m}.$
+Recall that the Fourier transform of convolution $\eta*\nu$ has the property $\widehat{\eta*\nu}(\zeta)=\hat{\eta}(\zeta)\hat{\nu}(\zeta),$ for details; see [25].
+It is said that $\eta$ is a Rajchman measure if $\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=0.$
 Definition 2.2.
@@
 Theorem 2.5.
-[14, Lemma 2.4.1, Lemma 2.4.2]
+[15, Lemma 2.4.1, Lemma 2.4.2]
   1. (1)
@@
     Let $E$ be a nonempty subset of $\mathbb{R}^{m},$ and $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ be a bi-Lipschitz map. Then, $\dim_{\mathcal{A}}T(E)=\dim_{\mathcal{A}}E.$
-Let $\mathcal{P}(\mathbb{R}^{m})$ be the set of probability measures on $\mathbb{R}^{m}$ with the metric defined as
-$d_{L}(\mu,\nu)=\sup_{Lip(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\mu-\int_{\mathbb{R}^{m}}fd\nu\Big{|}\Big{\}},$
-where $f$ is a Lipschitz function with the Lipschitz constant $Lip(f)$ equal to $1$, see [7, 25]. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, the convolution $\mu*\nu$ is given as
+Let $\mathcal{P}(\mathbb{R}^{m})$ be the set of Borel probability measures on $\mathbb{R}^{m}$ with the metric defined as
+$d_{L}(\mu,\nu)=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\mu-\int_{\mathbb{R}^{m}}fd\nu\Big|\Big\},$
+where $\text{Lip}_{1}$ is a set of all real-valued Lipschitz functions on $\mathbb{R}^{m}$ with the Lipschitz constant less than or equal to $1$, see [8, 30].
+Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, the convolution $\mu*\nu$ is given as
 $\mu*\nu(A)=\int_{\mathbb{R}^{m}}\mu(A-x)d\nu(x),\hskip 11.38092pt\mbox{for all Borel set }A\in\mathbb{R}^{m}.$
@@
 [14] studied the convolutions of equicontractive self-similar measures on $\mathbb{R}.$
+We now define $L^{q}$ dimension of a Borel probability measure with bounded support.
+This concept was initially introduced by Rényi [36], and is one of the fundamental ingredients in the study of multifractal formalism; see [22].
 Definition 2.6.
-[30] Let $q\in(1,\infty).$ If $\mu$ is a probability measure on $\mathbb{R}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
+[38] Let $q\in(1,\infty).$ If $\mu$ be a Borel probability measure on $\mathbb{R}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
 $\tau(\mu,q)=\liminf_{k\to\infty}\frac{-\log\sum_{I\in D_{k}}\mu(I)^{q}}{k},$
 where $D_{k}=\{[j2^{-k},(j+1)2^{-k}):j\in\mathbb{Z}\}$ is the level-$k$ dyadic partitions of $\mathbb{R}.$ The $L^{q}$ dimension of $\mu$ is given as $\dim_{L^{q}}(\mu)=\frac{\tau(\mu,q)}{q-1}.$
-If $\mu$ is a probability measure on $\mathbb{R}^{m}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
+If $\mu$ is a Borel probability measure on $\mathbb{R}^{m}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
 $\tau(\mu,q)=\liminf_{k\to\infty}\frac{-\log\sum_{I^{1}\times I^{2}\times\cdots\times I^{m}\in D_{k}^{m}}\mu(I^{1}\times I^{2}\times\cdots\times I^{m})^{q}}{k},$
@@
 For entropy and $L^{q}$ dimension, researchers consider preferring logarithms on base $2,$ for normalizing the value of dimension.
-Several recent works on $L^{q}$ dimension can be studied in [2, 24, 30].
+As $q\rightarrow 1^{+},$ the $L^{q}$ dimension tends to the Hausdorff dimension of the measure.
+Shmerkin [38] computed $L^{q}$ dimension of self-similar measure under ESC.
+Peres and Solomyak [31] computed the $L^{q}$ dimension of self-conformal measures, without any separation assumptions.
+Several recent works on $L^{q}$ dimension can be studied in [2, 27, 28, 29].
 Example 2.7.
-We now proceed to estimate the $L^{q}$ dimension of some well-known probability measures, to enhance clarity for new readers.
+We now proceed to estimate the $L^{q}$ dimension of some well-known Borel probability measures, to enhance clarity for new readers.
   • Let $x_{0}\in\mathbb{R}$ and $\delta_{x_{0}}$ be the Dirac measure supported on $x_{0}.$ Then,
@@
     This gives $\dim_{L^{q}}(\delta_{x_{0}})=0.$
-  • Let $c\in\mathbb{R}.$ Then,
-    $\tau(c\delta_{x_{0}},q)=\liminf_{k\rightarrow\infty}\frac{-\log(c)^{q}}{k}=0,\text{ and }\dim_{L^{q}}(c\delta_{x_{0}})=0.$
-  • Let $x_{1},x_{2},\ldots,x_{n}$ be distinct points in $\mathbb{R}.$ Then,
+  • Let $x_{1},x_{2},\ldots,x_{n}$ be distinct points in $\mathbb{R}$ and $c_{1},c_{2},\ldots,c_{n}>0.$ Then,
     $\tau(c_{1}\delta_{x_{1}}+c_{2}\delta_{x_{2}}+\dots+c_{n}\delta_{x_{n}},q)=\liminf_{k\rightarrow\infty}\frac{-\log(c_{1}^{q}+c_{2}^{q}+\ldots+c_{n}^{q})}{k}=0,$
@@
 (2.1) $\Delta_{n}=\inf\{d(h_{i},h_{j}):i,j\in\{1,2,\ldots,N\}^{n},i\neq j\}.$
-For $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots N\}^{n},$ we denote $h_{i}=h_{i_{1}}\circ h_{i_{2}}\circ\cdots\circ h_{i_{n}}.$ An IFS $\mathcal{I}$ is said to satisfy the exponential separation condition (ESC) if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ see [19].
-We would like to highlight that in some of the articles, the previous definition of ESC is mentioned with all arbitrarily large $n\in\mathbb{N}$ such that $\Delta_{n}>\delta^{n},$ while some others say that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N},$ for example refer to the literature [3, 18, 30].
-Observe that if $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ then both are true.
-In the other case, if $\Delta_{n}>\delta^{n},$ for all $n\geq n_{0},$ then choose $\delta_{*}=\min\{\delta,\Delta_{1},\sqrt{\Delta_{2}}\ldots,\sqrt[n_{0}]{\Delta_{n_{0}}}\}.$ This gives $\Delta_{n}>\delta_{*}^{n},$ for all $n\in\mathbb{N}.$ With reference to Theorem 1.5 of [17], for an IFS on $\mathbb{R}$ consisting of all the similarity mappings with the algebraic parameters, $\Delta_{n}$ is either zero or strictly greater than $\delta^{n},$ for some constant $\delta>0,$ establishing the equivalency among the three.
-We strongly believe that all three definitions are equivalent in general.
-Further, note that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N}$ is equivalent to $\Delta_{n}>c\delta^{n}$ for all $n\in\mathbb{N},$ for some constant $c\in\mathbb{R}.$ The direct part is straightforward.
-For the converse, consider two cases.
-First $c>1,$ then it is obvious.
-In the other case, assume $0<c\leq 1.$ Choose $\delta_{*}=c\delta,$ then $\Delta_{n}>c\delta^{n}>(c\delta)^{n}.$ An example was presented by Garsia [16] as $\{\mathbb{R}:~h_{1}(x)=\gamma^{-1}x+1,h_{2}(x)=\gamma^{-1}x-1\},$ where $\gamma$ is the real root of the equation $x^{3}-x^{2}-2=0.$ Hochman [18] shows that this IFS satisfies ESC, but not OSC.
+For $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots N\}^{n},$ we denote $h_{i}=h_{i_{1}}\circ h_{i_{2}}\circ\cdots\circ h_{i_{n}}.$ An IFS $\mathcal{I}$ is said to satisfy the exponential separation condition (ESC) if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n}$ for infinitely many $n\in\mathbb{N},$ where as $\mathcal{I}$ is said to satisfy the strong ESC if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N},$ see [3, 19, 20, 38].
+An example was presented by Garsia [17] as $\{\mathbb{R}:~h_{1}(x)=\gamma^{-1}x+1,h_{2}(x)=\gamma^{-1}x-1\},$ where $\gamma$ is the real root of the equation $x^{3}-x^{2}-2=0.$ Hochman [19] shows that this IFS satisfies ESC, but not OSC.
 Meanwhile, super-exponential condensation means $\lim_{n\rightarrow\infty}\frac{1}{n}\log\Delta_{n}=-\infty.$ Bárány and Käenmäki [3, Theorem 2.1] showed the existence of an uncountable family of homogeneous self-similar sets of $\mathbb{R}$ that do not have total overlaps and satisfy super-exponential condensation.
@@
 Let $\mathcal{I}_{1}=\{\mathbb{R}^{k};h_{1},h_{2},\ldots,h_{N}\}$ and $\mathcal{I}_{2}=\{\mathbb{R}^{m};g_{1},g_{2},\ldots,g_{N}\}$ be two IFSs such that $h_{i}(x)=cU_{i}x+a_{i}$ and $g_{i}(x)=cV_{i}x+b_{i},$ where $c\in(-1,1),a_{i}\in\mathbb{R}^{k},b_{i}\in\mathbb{R}^{m},U_{i}$ and $V_{i}$ are $k\times k$ and $m\times m$ orthogonal matrices, respectively.
-If $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, then the product IFS $\mathcal{I}_{1}\times\mathcal{I}_{2}$ given as
+If $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy strong ESC, then the product IFS $\mathcal{I}_{1}\times\mathcal{I}_{2}$ given as
 $\mathcal{I}_{1}\times\mathcal{I}_{2}=\{\mathbb{R}^{k+m};\psi_{ij}:i,j=1,2,\ldots,N\},\ \psi_{ij}=(h_{i},g_{j})$
-satisfies ESC.
+satisfies strong ESC.
 Proof.
-Since IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, there exist constants $\delta_{1}$ and $\delta_{2}>0,$ respectively such that
+Since IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy strong ESC, there exist constants $\delta_{1}$ and $\delta_{2}>0,$ respectively such that
 $\Delta_{n}^{\mathcal{I}_{1}}>\delta_{1}^{n}~~\text{ and }~~\Delta_{n}^{\mathcal{I}_{2}}>\delta_{2}^{n},$
@@
 $\geq\max\{\Delta_{n}^{\mathcal{I}_{1}},\Delta_{n}^{\mathcal{I}_{2}}\}>\min\{\delta_{1}^{n},\delta_{2}^{n}\}=\delta_{3}^{n},$
-where $\delta_{3}=\min\{\delta_{1},\delta_{2}\},$ for all $n\in\mathbb{N},$ assuring that $\mathcal{I}_{1}\times\mathcal{I}_{2}$ satisfies ESC. ∎
+where $\delta_{3}=\min\{\delta_{1},\delta_{2}\},$ for all $n\in\mathbb{N},$ assuring that $\mathcal{I}_{1}\times\mathcal{I}_{2}$ satisfies strong ESC. ∎
 Remark 3.3.
-The previous theorem plays a considerable role in the context of extending the Hochman [18, Theorem 1.4] groundbreaking result to the product of IFSs.
-Meanwhile, Bárány and Verma [5] introduced a weak ESC; it remains an open question whether the result holds in that case as well.
+The previous theorem plays a considerable role in the context of extending the Hochman [19, Theorem 1.4] groundbreaking result to the product of IFSs.
 Consider an IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ on $\mathbb{R}^{m}$ that consists of a finite number of affine mappings.
@@
 (3.1) $\Delta_{n}^{*}:=\inf\{\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A))):i,j\in\{1,2,\ldots,N\}^{n},i\neq j\}.$
-Then, IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ is said to satisfy the modified ESC if there exists a $\delta>0$ such that $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$
-Note 3.4.
-It is noted that several researchers attempted to generalise the concept of ESC introduced by Hochman [17, 18].
-In a way, Bárány and Verma [5] defined weak ESC for self-similar IFS, ESC for graph-directed self-similar IFS, ESC for common fixed point system.
+Then, IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ is said to satisfy the modified ESC if there exists a $\delta>0$ such that $\Delta_{n}^{*}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ Similarly, we can define modified strong ESC.
+It is worth emphasizing that our definition of ESC is natural from a geometric perspective and significantly convenient.
+Computability and applicability of ESC and modified ESC might be seen as one being more algebraic in nature and the other is more geometric in nature.
+It is noted that several researchers [4, 6, 35] attempt to generalize and weaken the concept of ESC introduced by Hochman [18, 19].
+In a way, Bárány and M.
+Verma [4] defined weak ESC for self-similar IFS, ESC for graph-directed self-similar IFS, ESC for common fixed point system.
 Above all that, the most fundamental advantage of the modified ESC is that it is defined not only for self-similar IFS but also extends to more general classes of IFS, including self-affine IFS and self-conformal IFS.
@@
 $\Delta_{n}^{*}=\inf\{\mathfrak{H}(h_{i}([0,1]),h_{j}([0,1])):i,j\in\{1,2\}^{n},i\neq j\}=\frac{2}{3^{n}}>\frac{1}{3^{n}}.$
+This clearly shows that the IFS satisfies both modified ESC and modified strong ESC.
 The second example can be considered as Sierpiński gasket on $\mathbb{R}^{2}.$ Its IFS is elaborated as $\{\mathbb{R}^{2}:~h_{1},h_{2},h_{3}\},$ where
@@
 $\Delta_{n}^{*}$ $=\inf\{\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle)):i,j\in\{1,2,3\}^{n},i\neq j\}.$
-Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2,3\}^{n}.$ Then, $h_{i}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(0,0),(\frac{1}{2^{n}},0),(\frac{1}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}})$ and $h_{j}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(\frac{1}{2^{n}},0),(\frac{1}{2^{n-1}},0),\\ (\frac{3}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}}).$ Then, $\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle))=\frac{1}{2^{n}}$ and $\Delta_{n}^{*}=\frac{1}{2^{n}}.$
+Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2,3\}^{n}.$ Then, $h_{i}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(0,0),(\frac{1}{2^{n}},0),(\frac{1}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}})$ and $h_{j}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(\frac{1}{2^{n}},0),(\frac{1}{2^{n-1}},0),\\ (\frac{3}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}}).$ Then, $\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle))=\frac{1}{2^{n}}$ and $\Delta_{n}^{*}=\frac{1}{2^{n}}.$ This clearly shows that the IFS satisfies both modified ESC and modified strong ESC.
 Remark 3.5.
@@
 Theorem 3.6.
-Consider an IFS consisting of similarity transformations $(h_{i}(x)=c_{i}U_{i}x+a_{i},i=1,2,\ldots,N)$ on $\mathbb{R}^{m},$ then there exists a constant $k\in\mathbb{R}$ such that $\Delta_{n}^{*}\leq k\Delta_{n}.$
+Consider an IFS consisting of similarity transformations $(h_{i}(x)=c_{i}U_{i}x+a_{i},i=1,2,\ldots,N)$ on $\mathbb{R}^{m},$ then there exists a constant $k_{*}\in\mathbb{R}$ such that $\Delta_{n}^{*}\leq k_{*}\Delta_{n},$ for all $n\in\mathbb{N}.$
 Proof.
@@
 $\inf_{y\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\leq\|h_{i}(x)-h_{j}(x)\|_{2}.$
-Taking supremum both side
+Taking the supremum of both sides, we get
 $\sup_{x\in\text{co}(A)}\inf_{y\in\text{co}(A)}\|h_{i}(x)-h_{j}(y)\|_{2}\leq\sup_{x\in\text{co}(A)}\|h_{i}(x)-h_{j}(x)\|_{2}.$
@@
 $\Delta_{n}^{*}\leq k_{*}\Delta_{n}.$
-This completes the proof.
+Clearly, this holds for all $n\in\mathbb{N}.$ This completes the proof.
 Proposition 3.7.
-Consider a homogeneous IFS consisting of similarity transformations $(h_{i}(x)=cx+d_{i},i=1,2,\ldots,N)$ on $\mathbb{R},$ then the above two definitions (mentioned in equations (2.1) and (3.1)) of ESC coincide, that is, $\Delta_{n}=\Delta_{n}^{*}.$
+Consider a homogeneous IFS consisting of similarity transformations $(h_{i}(x)=cx+d_{i},i=1,2,\ldots,N)$ on $\mathbb{R},$ then the above two definitions (mentioned in equations (2.1) and (3.1)) of ESC coincide, that is, $\Delta_{n}=\Delta_{n}^{*},$ for all $n\in\mathbb{N}.$
 Proof.
@@
 $\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))=|d_{i}-d_{j}|=d(h_{i},h_{j}).$
-Taking the infimum on both sides over $i,j\in\{1,2,\ldots,N\}^{n},$ we get $\Delta_{n}^{*}=\Delta_{n},$ concluding the proof.
-Note 3.9.
-If homogeneous IFS on $\mathbb{R}$ satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$ Thus, the homogeneous IFS satisfies modified ESC.
+Taking the infimum on both sides over $i,j\in\{1,2,\ldots,N\}^{n},$ we get $\Delta_{n}^{*}=\Delta_{n},$ for all $n\in\mathbb{N},$ concluding the proof.
+Note 3.8.
+If homogeneous IFS on $\mathbb{R}$ satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ Thus, the homogeneous IFS satisfies modified ESC.
 The next remark will give some hints on the ESC for the distinct contraction ratios.
@@
 $\frac{1}{\min\{b,1\}}\mathfrak{H}(h_{i}(\text{co}(A)),h_{j}(\text{co}(A)))\geq\{(c_{i}-c_{j})+(d_{i}-d_{j})\}\geq\frac{1}{k_{n}}d(h_{i},h_{j}).$
-Taking infimum over $i,j\in\{1,2,\ldots,N\}^{n}$ both sides, we get $\frac{1}{\min\{b,1\}}\Delta_{n}^{*}\geq\frac{1}{k_{n}}\Delta_{n}.$ If the IFS satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\frac{\min\{b,1\}}{k_{n}}\delta^{n}=\min\{b,1\}\Big{(}\min_{i\in\{1,2,\ldots,N\}}|c_{i}|\delta\Big{)}^{n},$ for all $n\in\mathbb{N}.$
-In the upcoming theorems, several dense subsets of $\mathcal{K}(\mathbb{R}^{m})$ are explored, which will be helpful in the dimension-preserving approximation of sets.
+Taking infimum over $i,j\in\{1,2,\ldots,N\}^{n}$ both sides, we get $\frac{1}{\min\{b,1\}}\Delta_{n}^{*}\geq\frac{1}{k_{n}}\Delta_{n}.$ If the IFS satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\frac{\min\{b,1\}}{k_{n}}\delta^{n}=\min\{b,1\}\Big(\min_{i\in\{1,2,\ldots,N\}}|c_{i}|\delta\Big)^{n},$ for infinitely many $n\in\mathbb{N},$ claiming that the IFS satisfies modified ESC.
+In the upcoming theorems, several dense subsets of $\chi(\mathbb{R}^{m})$ are explored, which will be helpful in the dimension-preserving approximation of sets.
 Theorem 3.10.
-For each $\alpha\in[0,m],$ the set $A_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{A}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
+For each $\alpha\in[0,m],$ the set $A_{\alpha}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{A}}E=\alpha\}$ is dense in $\chi(\mathbb{R}^{m}).$
 Proof.
-Let $\epsilon>0$ and $E\in\mathcal{K}(\mathbb{R}^{m}).$ Since $E$ is compact subset of $\mathbb{R}^{m},$ there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\bigcup_{j=1}^{n}B(e_{j},\epsilon),$ where $B(e_{j},\epsilon)$ is the open set centered at $e_{j}$ and radius $\epsilon.$ Subsequently, consider $F=\{e_{1},e_{2},\dots,e_{n}\}\in\mathcal{K}(\mathbb{R}^{m}).$ In view of definition of Hausdorff metric and Example 2.3, we get
+Let $\epsilon>0$ and $E\in\chi(\mathbb{R}^{m}).$ Since $E$ is compact subset of $\mathbb{R}^{m},$ there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\bigcup_{j=1}^{n}B(e_{j},\epsilon),$ where $B(e_{j},\epsilon)$ is the open set centered at $e_{j}$ and radius $\epsilon.$ Subsequently, consider $F=\{e_{1},e_{2},\dots,e_{n}\}\in\chi(\mathbb{R}^{m}).$ In view of definition of Hausdorff metric and Example 2.4, we get
 $\mathfrak{H}(E,F)<\epsilon~~\text{ and }~~\dim_{\mathcal{A}}F=0.$
-This shows that the assertion is true in the case of $\alpha=0.$ Again, assume $\alpha\in(0,m].$ Since the Assouad dimension satisfies the open set property (see Theorem 2.4), $\dim_{\mathcal{A}}B(e_{j},\epsilon)=m.$ Now using [8, Theorem 1], there exists a set $F_{e_{1}}\subset B(e_{1},\epsilon)$ such that $\dim_{\mathcal{A}}F_{e_{1}}=\alpha.$ Define $F^{\prime}=F\cup F_{e_{1}}.$ In view of the finite stability of the Assouad dimension, we get
+This shows that the assertion is true in the case of $\alpha=0.$ Again, assume $\alpha\in(0,m].$ Since the Assouad dimension satisfies the open set property (see Theorem 2.5), $\dim_{\mathcal{A}}B(e_{j},\epsilon)=m.$ Now using [9, Theorem 1], there exists a set $F_{e_{1}}\subset B(e_{1},\epsilon)$ such that $\dim_{\mathcal{A}}F_{e_{1}}=\alpha.$ Define $F^{\prime}=F\cup F_{e_{1}}.$ In view of the finite stability of the Assouad dimension, we get
 $\dim_{\mathcal{A}}F^{\prime}=\max\{\dim_{\mathcal{A}}F,\dim_{\mathcal{A}}F_{e_{1}}\}=\alpha,$
@@
 Theorem 3.11.
-For each $\alpha\in[0,m],$ the set $H_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
+For each $\alpha\in[0,m],$ the set $H_{\alpha}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{H}}E=\alpha\}$ is dense in $\chi(\mathbb{R}^{m}).$
 Proof.
@@
 Theorem 3.12.
-The set $\mathcal{D}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E\neq\dim_{\mathcal{A}}E\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
+The set $\mathcal{D}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{H}}E\neq\dim_{\mathcal{A}}E\}$ is dense in $\chi(\mathbb{R}^{m}).$
 Proof.
-In case of $m=1,$ using [14, Theorem 2.1.1], one can clearly show that $\mathcal{D}$ is a non-empty subset of $\mathcal{K}(\mathbb{R}^{m}).$ The Bedford-McMullen carpet (denoted as $\mathcal{B}\subset\mathbb{R}^{2}$) has the Hausdorff dimension strictly less than the Assouad dimension (see [15, Theorem 2.1]) is an example in case of $m=2.$ For a higher dimension $(m>2)$, consider the set $\mathcal{B}\times\{(\underset{m-2\text{ times}}{\underbrace{0,0,\dots,0}})\}=\{(x,y,0,0,\ldots,0)\in\mathbb{R}^{m}:(x,y)\in\mathcal{B}\},$ concluding the set $\mathcal{D}$ is non-empty.
-Further, assume $E\in\mathcal{K}(\mathbb{R}^{m})$ and $\epsilon>0.$ Then, there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\cup_{j=1}^{n}B(e_{j},\epsilon).$ Observe that the Hausdorff and the Assouad dimensions are invariant under scaling and translation (see [11, 14]).
+In case of $m=1,$ using [15, Theorem 2.1.1], one can clearly show that $\mathcal{D}$ is a non-empty subset of $\chi(\mathbb{R}^{m}).$ The Bedford-McMullen carpet (denoted as $\mathcal{B}\subset\mathbb{R}^{2}$) has the Hausdorff dimension strictly less than the Assouad dimension (see [16, Theorem 2.1]) is an example in case of $m=2.$ For a higher dimension $(m>2)$, consider the set $\mathcal{B}\times\{(\underset{m-2\text{ times}}{\underbrace{0,0,\dots,0}})\}=\{(x,y,0,0,\ldots,0)\in\mathbb{R}^{m}:(x,y)\in\mathcal{B}\},$ concluding the set $\mathcal{D}$ is non-empty.
+Further, assume $E\in\chi(\mathbb{R}^{m})$ and $\epsilon>0.$ Then, there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\cup_{j=1}^{n}B(e_{j},\epsilon).$ Observe that the Hausdorff and the Assouad dimensions are invariant under scaling and translation (see [12, 15]).
 Using translation and scaling, we get $\mathcal{B}\times\{(0,0,\ldots,0)\}\subset B(e_{1},\epsilon).$ Choose $F=\{e_{1},e_{2},\ldots,e_{n}\}\cup(\mathcal{B}\times\{(0,0,\ldots,0)\}).$ Since the Assouad and the Hausdorff dimensions satisfy finite stability, we get
@@
   • A linear subspace $V\subseteq\mathbb{R}^{m}$ is said to be $D\mathcal{I}$-invariant if it is invariant under the orthogonal parts, i.e., $U_{i}(V)\subseteq V,$ for all $i.$
-We now mention the following theorem, which is a conclusion of [18, Theorem 1.4].
+We now mention the following theorem, which is a conclusion of [19, Theorem 1.4].
 Theorem 3.13.
@@
 It is noted that the IFS does not satisfy SSC, but there exists a subcollection of $2$-nd level IFS satisfying SSC (see [18, 43]).
-In view of this, the previous theorem is a weaker version of the Hochman result [18, Theorem 1.4].
-To our knowledge, we could not find any proof or statement of the following lemma in the literature; however, some partial mentions for Dirac measures can be seen in [30] and also, since convolution is the smoothing operation, we believe $\dim_{L^{q}}(\mu*\nu)\geq\dim_{L^{q}}(\mu).$ So, it is crucial to include the complete proof with all details for the reader’s convenience.
+In view of this, the previous theorem is a weaker version of the Hochman result [19, Theorem 1.4].
+The upcoming lemmas are elementary observations on the behaviour of the $L^{q}$ dimension, and their proofs are accommodated for young readers.
+It is remarkable here that Feng et al.
+[14, Lemma 2.2(ii)] proved one-sided inequality for general Borel probability measures by defining $L^{q}$ dimension in terms of integration.
 Lemma 3.16.
-Let $\mathcal{S}(\mathbb{R}^{m})=\big{\{}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}:x_{j}\in\mathbb{R}^{m},c_{j}>0,\sum_{j=1}^{n}c_{j}=1,n\in\mathbb{N}\big{\}}$ and $\mu$ be a probability measure with finite support.
+Let $\mathcal{S}(\mathbb{R}^{m})=\big\{\sum_{j=1}^{n}c_{j}\delta_{x_{j}}:x_{j}\in\mathbb{R}^{m},c_{j}>0,\sum_{j=1}^{n}c_{j}=1,n\in\mathbb{N}\big\}$ and $\mu$ be a Borel probability measure with finite support.
 Then, for a fixed $\omega\in\mathcal{S}(\mathbb{R}^{m}),$
@@
 Following the similar approach, we get $\dim_{L^{q}}(\mu)\geq\dim_{L^{q}}(\nu).$ This completes the proof.
+The upcoming theorems interestingly explore the approximation aspects of a Borel probability measure with a sequence of Borel probability measures preserving
+  • $L^{q}$ dimension.
+  • Rajchman measures.
+  • $L^{q}$ dimension and Rajchman.
+  • $L^{q}$ dimension and Fourier transform with power Fourier decay.
+Readers obviously think that, why we are interested in highlighting and studying some specific dense subsets of the space of Borel probability measures concerning the $L^{q}$ dimension and Rajchman or power Fourier decay.
+The Rajchman character of a measure essentially provides information about local regularity and is closely connected to the dimension theory.
+The Riemann-Lebesgue lemma says that a measure $\eta\in\mathcal{P}(\mathbb{R}^{m})$ is Rajchman whenever it is absolutely continuous with respect to the Lebesgue measure; refer to [33], and the references cited therein.
+Thus, if $\eta$ is not Rajchman, then it is singular.
+As an example, the Bernoulli convolutions, except for the particular set of parameters, satisfy the Rajchman property.
+If for some constant $C,$ we have $|\hat{\eta}(x)|\leq C|x|^{-\frac{t}{2}},$ for all $x\in\mathbb{R}^{m},$ then any set which supports $\eta$ has the Hausdorff dimension atleast $t.$ In fractal geometry, the classification or identification of a class of measures or IFS satisfying specific properties is always of great interest to researchers.
+Recently, Bárány and Käenmäki [3] showed that self-similar sets on $\mathbb{R}$ which neither satisfy ESC nor have exact overlaps are uncountable.
+Olsen [27] constructed a $G_{\delta}$ dense set in the space of the Borel probability measure on a compact subset of $\mathbb{R}^{m}$ with upper $L^{q}$ dimension equal to zero with respect to metric $d_{L}$.
+Recent researches clearly indicate that we want to explore the space $\mathcal{P}(\mathbb{R}^{m})$ and are curious to know whether almost (or typically) all Borel probability measures are Rajchman, or all self-affine measures are Rajchman, and so on.
+Rapaport [33] commented on the Rajchman property of self-similar measures on $\mathbb{R}^{m}.$ So, going through all the recent works, we thought, can we get some dense subsets of $\mathcal{P}(\mathbb{R}^{m})$ preserving $L^{q}$ dimension or Rajchman property?
+Our upcoming theorems answer these queries interestingly.
+Now, the future remarks what we can comment on all these upcoming dense sets to be $G_{\delta}$ dense.
 Theorem 3.18.
@@
 $d_{L}(\mu,\nu)<\frac{\epsilon}{2}.$
-Further, we consider a non-zero compactly supported measure $\lambda\in\mathcal{S}_{\beta}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Lemmas 3.17 and 3.18 implies that $\dim_{L^{q}}(\lambda^{\prime})=\dim_{L^{q}}(\lambda_{1})=\dim_{L^{q}}(\lambda)=\beta,$ that is, $\lambda^{\prime}\in\mathcal{S}_{\beta}.$ In view of the properties of convolution mentioned in Section 2, we get
-$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\nu-\int_{\mathbb{R}^{m}}fd\lambda^{\prime}\Big{|}\Big{\}}$
-$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\Big{(}\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big{)}-\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})\Big{|}\Big{\}}.$
+Further, we consider a non-zero compactly supported measure $\eta\in\mathcal{S}_{\beta}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\eta^{\prime}=\nu*\eta_{1},$ where $\eta_{1}(D):=\eta(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Lemmas 3.16 and 3.17 implies that $\dim_{L^{q}}(\eta^{\prime})=\dim_{L^{q}}(\eta_{1})=\dim_{L^{q}}(\eta)=\beta,$ that is, $\eta^{\prime}\in\mathcal{S}_{\beta}.$ In view of the properties of convolution mentioned in Section 2, we get
+$d_{L}(\nu,\eta^{\prime})$ $=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\nu-\int_{\mathbb{R}^{m}}fd\eta^{\prime}\Big|\Big\}$
+$=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\Big(\sum_{j=1}^{n}c_{j}\delta_{x_{j}}\Big)-\int_{\mathbb{R}^{m}}fd(\nu*\eta_{1})\Big|\Big\}.$
 Notice that
-$\int_{\mathbb{R}^{m}}fd(\nu*\lambda_{1})$ $=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\nu(x)d\lambda_{1}(y)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d(\sum_{j=1}^{n}c_{j}\delta_{x_{j}})(x)d\lambda_{1}(y)$
-$=\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j}+y)d\lambda_{1}(y).$
+$\int_{\mathbb{R}^{m}}fd(\nu*\eta_{1})$ $=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d\nu(x)d\eta_{1}(y)=\int_{\mathbb{R}^{m}}\int_{\mathbb{R}^{m}}f(x+y)d(\sum_{j=1}^{n}c_{j}\delta_{x_{j}})(x)d\eta_{1}(y)$
+$=\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j}+y)d\eta_{1}(y).$
 Thus,
-$d_{L}(\nu,\lambda^{\prime})$ $=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}f(x_{j})-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}\hskip 11.38092pt$
-$=\sup_{\text{Lip}(f)\leq 1}\Big{\{}\Big{|}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j})d\lambda_{1}(y)-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\lambda_{1}(y)\Big{|}\Big{\}}$
-$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}|f(x_{j})-f(y+x_{j})|d\lambda_{1}(y)\Big{\}}$
-$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}.$
-We claim that the support of $\lambda_{1}$ is a subset of $\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ Let $C\subset\mathbb{R}^{m}\setminus\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}.$ If $x\in C,$ then $\|x\|_{2}>\frac{\epsilon}{4\sqrt{m}}$ and $\frac{8\sqrt{m}diam(K)\|x\|_{2}}{\epsilon}>2diam(K).$ So, $\lambda_{1}(C)=\lambda(\frac{8\sqrt{m}diam(K)C}{\epsilon})=0.$ The previous estimation is reduced to
-$d_{L}(\nu,\lambda^{\prime})$ $\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\|y\|_{2}d\lambda_{1}(y)\Big{\}}$
-$\leq\sup_{\text{Lip}(f)\leq 1}\Big{\{}\int_{\big{[}\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big{]}^{m}}\text{Lip}(f)\frac{\epsilon}{2}d\lambda_{1}(y)\Big{\}}\leq\frac{\epsilon}{2}.$
+$d_{L}(\nu,\eta^{\prime})$ $=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\sum_{j=1}^{n}c_{j}f(x_{j})-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\eta_{1}(y)\Big|\Big\}\hskip 11.38092pt$
+$=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(x_{j})d\eta_{1}(y)-\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}f(y+x_{j})d\eta_{1}(y)\Big|\Big\}$
+$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}|f(x_{j})-f(y+x_{j})|d\eta_{1}(y)\Big\}$
+$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\sum_{j=1}^{n}c_{j}\int_{\mathbb{R}^{m}}\text{Lip}(f)\|y\|_{2}d\eta_{1}(y)\Big\}.$
+We claim that the support of $\eta_{1}$ is a subset of $\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}.$ Let $C\subset\mathbb{R}^{m}\setminus\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}.$ If $x\in C,$ then $\|x\|_{2}>\frac{\epsilon}{4\sqrt{m}}$ and $\frac{8\sqrt{m}diam(K)\|x\|_{2}}{\epsilon}>2diam(K).$ So, $\eta_{1}(C)=\eta(\frac{8\sqrt{m}diam(K)C}{\epsilon})=0.$ The previous estimation is reduced to
+$d_{L}(\nu,\eta^{\prime})$ $\leq\sup_{f\in\text{Lip}_{1}}\Big\{\int_{\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}}\text{Lip}(f)\|y\|_{2}d\eta_{1}(y)\Big\}$
+$\leq\sup_{f\in\text{Lip}_{1}}\Big\{\int_{\big[\frac{-\epsilon}{4\sqrt{m}},\frac{\epsilon}{4\sqrt{m}}\big]^{m}}\text{Lip}(f)\frac{\epsilon}{2}d\eta_{1}(y)\Big\}\leq\frac{\epsilon}{2}.$
 Using the above estimate, we immediately get
-$d_{L}(\mu,\lambda^{\prime})\leq d_{L}(\mu,\nu)+d_{L}(\nu,\lambda^{\prime})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon,$
+$d_{L}(\mu,\eta^{\prime})\leq d_{L}(\mu,\nu)+d_{L}(\nu,\eta^{\prime})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon,$
 showing that $\mathcal{S}_{\beta}$ is dense.
-Definition 3.20.
-[22] The Fourier transform of $\mu\in\mathcal{P}(\mathbb{R}^{m})$ is defined as
-$\hat{\mu}(\xi)=\int e^{-2\pi i\xi.x}d\mu(x),~~~~\xi\in\mathbb{R}^{m}.$
-Recall that the Fourier transform of convolution $\mu*\nu$ has the property $\widehat{\mu*\nu}(\xi)=\hat{\mu}(\xi)\hat{\nu}(\xi),$ for details; see [22].
-It is said that $\mu$ is a Rajchman measure if $\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$ The Riemann-Lebesgue lemma says that $\mu$ is Rajchman whenever it is absolutely continuous with respect to the Lebesgue measure; refer to [27], and the references cited therein.
-Solomyak [32, 33] studied the Fourier decay for self-similar measures and homogeneous self-affine measures.
 Proposition 3.19.
-If the Fourier transform of a measure $\nu$ is bounded and $\mu$ is Rajchman (with power Fourier decay), then $\mu*\nu$ is also a Rajchman (with power Fourier decay).
-For any Rajchman measure $\mu$ and real number $\beta>0$, we have $\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=0$, where $\nu(D)=\mu(\beta D)$ for each Borel set $D$.
+  • (i)
+    Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ If $\mu$ is Rajchman (with power Fourier decay), then $\mu*\nu$ is also a Rajchman (with power Fourier decay).
+  • (ii)
+    Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ be any Rajchman measure. Let $\beta>0$ be any real number and $\eta(D)=\mu(\beta D)$ for each Borel set $D$. Then,
+    $\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=0.$
 Proof.
-Since $\nu$ is bounded, $\nu(\mathbb{R}^{m})<\infty.$ Consider
-$\lim_{|\xi|\to\infty}|\widehat{\mu*\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)||\hat{\nu}(\xi)|=K\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0.$
+It is well-known (and obvious) that the Fourier transform of a finite measure is bounded.
+As $\nu\in\mathcal{P}(\mathbb{R}^{m}),$ there exists $K\in\mathbb{R}$ such that $|\hat{\nu}(\zeta)|\leq K,$ for all $\zeta\in\mathbb{R}^{m}.$ Consider
+$\lim_{|\zeta|\to\infty}|\widehat{\mu*\nu}(\zeta)|=\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)\hat{\nu}(\zeta)|=\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)||\hat{\nu}(\zeta)|\leq K\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0.$
 Again, note that
-$\lim_{|\xi|\to\infty}|\hat{\nu}(\xi)|=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.x}d\nu(x)\Big{|}=\lim_{|\xi|\to\infty}\Big{|}\int e^{-2\pi i\xi.\beta x}d\mu(x)\Big{|}=0,$
+$\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=\lim_{|\zeta|\to\infty}\Big|\int e^{-2\pi i\zeta.x}d\eta(x)\Big|=\lim_{|\zeta|\to\infty}\Big|\int e^{-2\pi i\zeta.\beta x}d\mu(x)\Big|=0,$
 as $\mu$ is Rajchman, concluding the assertion. ∎
 Theorem 3.20.
-The set $\mathcal{S}_{R}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$.
-Moreover, for each $\beta\in[0,m]$, the sets $\mathcal{S}_{R}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\lim_{|\xi|\to\infty}|\hat{\mu}(\xi)|=0\}$ and $\mathcal{S}_{F}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\hat{\mu}(\xi)\text{ has power Fourier decay}\}$ are dense in $\mathcal{P}(\mathbb{R}^{m})$.
+The set $\mathcal{S}_{R}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0\}$ is dense in $\mathcal{P}(\mathbb{R}^{m})$.
+Moreover, for each $\beta\in[0,m]$, the sets $\mathcal{S}_{R}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\lim_{|\zeta|\to\infty}|\hat{\mu}(\zeta)|=0\}$ and $\mathcal{S}_{F}^{\beta}:=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\dim_{L^{q}}(\mu)=\beta\text{ and }\hat{\mu}(\zeta)\text{ has power Fourier decay}\}$ are dense in $\mathcal{P}(\mathbb{R}^{m})$.
 Proof.
@@
 $d_{L}(\mu,\nu)<\frac{\epsilon}{2}.$
-Choose a non-zero compactly supported Rajchman measure $\lambda\in\mathcal{S}_{R}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\lambda^{\prime}=\nu*\lambda_{1},$ where $\lambda_{1}(D):=\lambda(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Proposition 3.21 implies that $\lambda^{\prime}\in\mathcal{S}_{R},$ that is, $\lambda^{\prime}$ is Rajchman.
-Following the lines of Theorem 3.19, one gets $d_{L}(\mu,\lambda^{\prime})<\epsilon,$ showing that $\mathcal{S}_{R}$ is dense in $\mathcal{P}(\mathbb{R}^{m}).$ Further, to show $\mathcal{S}_{R}^{\beta}$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ choose $\lambda\in\mathcal{S}_{R}^{\beta}.$ This concludes the proof.
+Choose a non-zero compactly supported Rajchman measure $\eta\in\mathcal{S}_{R}$ with its support $K\subset\mathbb{R}^{m}.$ Let $\eta^{\prime}=\nu*\eta_{1},$ where $\eta_{1}(D):=\eta(\frac{8\sqrt{m}diam(K)D}{\epsilon})$ and $diam(K):=\sup\{\|x-y\|_{2}:x,y\in K\}.$ Note that Proposition 3.19 implies that $\eta^{\prime}\in\mathcal{S}_{R},$ that is, $\eta^{\prime}$ is Rajchman.
+Following the lines of Theorem 3.18, one gets $d_{L}(\mu,\eta^{\prime})<\epsilon,$ showing that $\mathcal{S}_{R}$ is dense in $\mathcal{P}(\mathbb{R}^{m}).$ Further, to show $\mathcal{S}_{R}^{\beta}$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ choose $\eta\in\mathcal{S}_{R}^{\beta}.$ This concludes the proof.
 ∎
@@
 […]
-Theorem 4.1.
-Let $0\leq\alpha\leq m$ and $\mu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, there exist Borel probability measures $\nu,\lambda$ such that $\mu=\nu*\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$
-Theorem 4.2.
-Let $0\leq\alpha\leq m$, and $\mu$ be a finite Borel measure such that $\dim_{L^{q}}(\mu)\leq\alpha.$ Then, there exist finite Borel measures $\nu,\lambda$ such that $\mu=\nu+\lambda$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\lambda)=\alpha.$
+Issue 1.
+Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ and $\alpha\in\mathbb{R}$ with $0\leq\alpha\leq\dim_{L^{q}}(\mu).$ Then, there exist Borel probability measures $\nu,\eta$ such that $\mu=\nu*\eta$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\eta)=\alpha.$
+Issue 2.
+Let $0\leq\alpha\leq m$, and $\mu$ be a finite Borel measure such that $\dim_{L^{q}}(\mu)\leq\alpha.$ Then, there exist finite Borel measures $\nu,\eta$ such that $\mu=\nu+\eta$ and $\dim_{L^{q}}(\nu)=\dim_{L^{q}}(\eta)=\alpha.$
 Statements and Declarations